### v2

Abstract
Bregman proximal-type algorithms (BPs), such as mirror descent, have become popular tools in machine learning and data science for exploiting problem structures through non-Euclidean geometries. In this paper, we show that BPs can get trapped near a class of non-stationary points, which we term spurious stationary points. Such stagnation can persist for any finite number of iterations if the gradient of the Bregman kernel is not Lipschitz continuous, even in convex problems. The root cause lies in a fundamental contrast in descent behavior between Euclidean and Bregman geometries: While Euclidean gradient descent ensures sufficient decrease near any non-stationary point, BPs may exhibit arbitrarily slow decrease around spurious stationary points. As a result, commonly used Bregman-based stationarity measure, such as relative change in terms of Bregman divergence, can vanish near spurious stationary points. This may misleadingly suggest convergence, even when the iterates remain far from any true stationary point. Our analysis further reveals that spurious stationary points are not pathological, but rather occur generically in a broad class of nonconvex problems with polyhedral constraints. Taken together, our findings reveal a serious blind spot in Bregman-based optimization methods and calls for new theoretical tools and algorithmic safeguards to ensure reliable convergence.
1 Introduction
In this paper, we consider structured nonsmooth (non)convex optimization problems of the form
$\min\limits_{\bm{x}\in{\mathbb{R}}^{n}}F(\bm{x}):=f(\bm{x})+g(\bm{x}),$ (P)
where $\operatorname{dom}(g)={\cal X}$ is a nonempty closed convex set, $f:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$ is a continuously differentiable function, and $g:{\mathbb{R}}^{n}\rightarrow\overline{\mathbb{R}}$ is a convex and locally Lipschitz continuous function.
To solve (P), Bregman proximal-type algorithms (BPs) are widely used for leveraging the geometry of ${\cal X}$ while avoiding costly Euclidean projections or proximal operations; see, e.g., (Beck and Teboulle, 2003; Birnbaum et al., 2011; Arora et al., 2012; Zhang et al., 2021).
In this work, we study BPs under a unified update rule of the form:
$\bm{x}^{k+1}=\mathop{\operatorname*{argmin}}\limits_{\bm{y}\in{\mathbb{R}}^{n}}\ \left\{\gamma\left(\bm{y};\bm{x}^{k}\right)+g(\bm{y})+\frac{1}{t_{k}}D_{h}(\bm{y},\bm{x}^{k})\right\},$ (1)
where $\gamma(\ \cdot\ ;\bm{x})$ is the surrogate model for $f$ at point $\bm{x}$, $t_{k}\geq 0$ is the step size, and $D_{h}$ denotes the Bregman divergence induced by a kernel function $h$. Many classical algorithms fit within this framework. For example, setting $\gamma(\bm{y};\bm{x}^{k})=f(\bm{x}^{k})+\nabla f(\bm{x}^{k})^{T}(\bm{y}-\bm{x}^{k})$ recovers the Bregman proximal gradient method (BPG) (Censor and Zenios, 1992; Bauschke et al., 2017, 2019; Zhu et al., 2021). Choosing $\gamma=f$ gives the Bregman proximal point method (Chen and Teboulle, 1993; Kiwiel, 1997). Moreover, using a second-order surrogate, $\gamma(\bm{y};\bm{x}^{k}):=f(\bm{x}^{k})+\nabla f(\bm{x}^{k})^{T}(\bm{y}-\bm{x}^{k})+\frac{1}{2}(\bm{y}-\bm{x}^{k})^{T}\nabla^{2}f(\bm{x}^{k})(\bm{y}-\bm{x}^{k})$, leads to a second-order variant recently studied by Doikov and Nesterov (2023).
The central finding of this paper is that BPs can become trapped near certain non-stationary points, which we term spurious stationary points, and fail to escape within any finite number of iterations when the gradient of the Bregman kernel $h$ is not Lipschitz continuous. This behavior stands in sharp contrast to Euclidean gradient methods, as well as to BPs equipped with either full-domain kernels (i.e., $\operatorname{dom}(h)={\mathbb{R}}^{n}$) (Zhang, 2024) or kernels with Lipschitz continuous gradients (Zhang and He, 2018). These regularity conditions ensure the mirror map is invertible and well-conditioned, under which Bregman geometry aligns with its Euclidean counterpart.
To illustrate the failure mode in the absence of such regularity, which is often the case in practical applications where Bregman and Euclidean geometries are misaligned, we next present a simple linear programming (LP) problem where BPG becomes trapped near a non-optimal solution.
Illustrative Example – Pathological Behavior of BPG. We consider the following simple LP problem:
$\begin{array}[]{rl}\min\limits_{x_{1},x_{2}}&-x_{1}\\ {\rm s.t.}&x_{1}+x_{2}=1,x_{1},x_{2}\geq 0,\end{array}$ (c-ex)
which admits the unique solution at $(1,0)$. If we choose the Boltzmann–Shannon entropy kernel, i.e., $h(\bm{x})=\sum_{i=1}^{2}x_{i}\log x_{i}$, BPG admits the closed-form iteration:
$\bm{x}^{k+1}$ $=\left(\frac{x_{1}^{k}}{x_{1}^{k}+e^{-t}x_{2}^{k}},\frac{e^{-t}x_{2}^{k}}{x_{1}^{k}+e^{-t}x_{2}^{k}}\right),\quad\forall~{}k\in{\mathbb{N}}_{+}.$
A key observation is that, for any finite number of iterations $k\leq K$ and any tolerance $\epsilon>0$, one can construct a feasible initialization such that every iterate $\bm{x}_{k}$ remains within an $\epsilon$-neighborhood of the spurious stationary point $(0,1)$, a non-optimal feasible point where BPG stagnates. See the trajectory plot below.
Figure 1: The trajectory of BPG with the Boltzmann–Shannon entropy kernel on the LP instance  (c-ex). For a suitably chosen initialization (possibly depending on $K$ and $\epsilon$), all iterates $\{x_{k}\}_{k\in[K]}$ remain trapped near the non-optimal point $(0,1)$, exhibiting finite-time stagnation. The shaded region indicates the predefined $\epsilon$-neighborhood.
The above example prompts a natural question: How can we rigorously define spurious stationary points and characterize the finite-time stagnation behavior of BPs near them? In what follows, we formalize this phenomenon and present our main theoretical results.
We begin by precisely characterizing the commonly used Bregman-based stationarity measure, such as $D_{h}(\bm{x}^{k+1},\bm{x}^{k})$, where the surrogate model $\gamma$ may vary as discussed above. In Theorem 1, we show that for any feasible sequence $\{\bm{z}^{k}\}_{k\geq 0}$ converging to a point $\overline{\bm{z}}$, the vanishing of the stationarity measure is equivalent to the existence of a limit subgradient at $\overline{\bm{z}}$, whose entries are zero for all coordinates $i$ where $\overline{z}_{i}\in{\rm int}(\operatorname{dom}(h))$. This equivalence naturally motivates our definition of spurious stationary points, see Definition 3. Notably, the occurrence of such points depends solely on the choice of the Bregman kernel. For full-domain kernels (i.e., $\operatorname{dom}(h)={\mathbb{R}}^{n}$), the interior includes all coordinates, thereby eliminating spurious stationary points entirely.
Building on this characterization, we prove a computational hardness result for BPs, i.e., Theorem 2: For any fixed iteration budget $K$, there exists a feasible initialization sufficiently close to a spurious stationary point such that the iterates remain trapped within its neighborhood for all $k\leq K$. This result formalizes the finite-time stagnation behavior of BPs and underscores their inability to escape from spurious stationary points under Bregman geometry.
To understand how broadly this issue arises, we further show in Theorem 3 that spurious stationary points are structurally ubiquitous in constrained nonconvex optimization. Specifically, for any kernel with domain $\operatorname{dom}(h)={\mathbb{R}}_{+}^{n}$, every vertex of a polyhedral constraint set that is not a true stationary point becomes a spurious stationary point under Bregman geometry.
Finally, although Theorem 2 guarantees finite-time stagnation for BPs initialized near a spurious stationary point located on the boundary of $\operatorname{dom}(h)$, our findings go further. In Section 4, we construct a nonconvex problem (see Example 3) where the iterates of BPs become trapped near a spurious stationary point, even when initialized at a well-behaved interior point located far from any spurious stationary point. This striking example shows that this finite-time stagnation phenomemon is not merely a consequence of poor initialization but a fundamental algorithmic vulnerability: BPs can be drawn toward spurious stationary points during their course of iteration and fail to recover.
Overall, our results reveal a critical flaw in existing Bregman proximal-type algorithms: For nonconvex problems, finite-time convergence to an approximately stationary point cannot be reliably guaranteed, even when the stationarity measures are small. This highlights the need for new theoretical tools and algorithmic strategies that can better cope with the challenges under Bregman geometry.
Notation.
We denote by $\overline{{\mathbb{R}}},\,{\mathbb{R}},\,{\mathbb{R}}_{+}$ the sets of extended real numbers, real numbers, and nonnegative real numbers, respectively. For a vector $\bm{x}\in{\mathbb{R}}^{n}$, its $i$-th coordinate is represented by $x_{i}$, and $\bm{x}_{\mathcal{I}}$ denotes a subvector of $\bm{x}$ indexed by $\mathcal{I}$. The Euclidean ball ${\mathbb{B}}_{\epsilon}(\bm{x})$ is defined as ${\mathbb{B}}_{\epsilon}(\bm{x}):=\{\bm{y}\in{\mathbb{R}}^{n}:\|\bm{x}-\bm{y}\|_{2}\leq\epsilon\}$. Given a set $\mathcal{X}\subseteq{\mathbb{R}}^{n}$, we use $\operatorname{cl}(\mathcal{X})$, ${\rm int}(\mathcal{X})$, and $\operatorname{bd}(\mathcal{X})$ to denote its closure, interior, and boundary, respectively. The indicator function $\delta_{\mathcal{X}}$ of a set $\mathcal{X}$ is defined as $\delta_{\mathcal{X}}(\bm{x})=0$ if $\bm{x}\in\mathcal{X}$; $\delta_{\mathcal{X}}(\bm{x})=+\infty$ otherwise. We employ the shorthand $\delta_{\bm{g}(\bm{x})={\mathbf{0}}}$ to compactly represent $\delta_{\{\bm{x}\in{\mathbb{R}}^{n}:\bm{g}(\bm{x})={\mathbf{0}}\}}$ for any real-valued function $\bm{g}:{\mathbb{R}}^{n}\rightarrow{\mathbb{R}}$. Unless otherwise specified, the sequence $\{\bm{x}^{k}\}_{k\geq 0}$ always refers to the iterates generated by BPs. In contrast, the sequences $\{\bm{y}^{k}\}_{k\geq 0}$ and $\{\bm{z}^{k}\}_{k\geq 0}$ are auxiliary sequences introduced solely for technical analysis.
Organization.
The remainder of the paper is organized as follows. Section 2 introduces the problem setup and necessary preliminaries. In Section 3, we present our three main theoretical results. Section 4 provides two illustrative examples that demonstrate the practical implications of the computational hardness of BPs. The proofs of our main results are given in Section 5. We conclude with final remarks in Section 6.
2 Preliminaries and Problem Setup
In this section, we introduce the key assumptions and definitions that form the foundation of our analysis. We begin with the definition of a separable kernel function.
Definition 1.
A function $h:{\mathbb{R}}^{n}\rightarrow\overline{\mathbb{R}}$ is called a separable kernel function if it satisfies the following conditions:
  1. (i)
    There exists a univariate function $\varphi:{\mathbb{R}}\rightarrow\overline{\mathbb{R}}$ such that $h(\bm{x})=\sum_{i=1}^{n}\varphi(x_{i}),$ where $\varphi$ is continuously differentiable on ${\rm{int}}(\operatorname{dom}(\varphi))$.
  2. (ii)
    For every sequence $\{x^{k}\}_{k\geq 0}\subset\mathrm{int}(\operatorname{dom}(\varphi))$ converging to a point $x\in\mathrm{bd}(\operatorname{dom}(\varphi))$, we have $|\varphi^{\prime}(x^{k})|\to+\infty$.
  3. (iii)
    The function $\varphi$ is strictly convex.
The separability structure in property (i) is prevalent in a broad range of applications; see, e.g., (Bauschke et al., 2019; Azizian et al., 2022; Li et al., 2023). Properties (ii) and (iii) are collectively referred to as Legendre-type conditions, as introduced in Rockafellar (1970, Chapter 26). The following are common examples of kernel functions that satisfy Definition 1:
Example 1.
(See (Bauschke et al., 2017, Example 1).)
  1. (i)
    Boltzmann–Shannon entropy kernel $h(\bm{x})=\sum_{i=1}^{n}x_{i}\log(x_{i})$;
  2. (ii)
    Fermi–Dirac entropy kernel $h(\bm{x})=\sum_{i=1}^{n}x_{i}\log(x_{i})+(1-x_{i})\log(1-x_{i})$;
  3. (iii)
    Burg entropy kernel $h(\bm{x})=\sum_{i=1}^{n}-\log(x_{i})$;
  4. (iv)
    Fractional power kernel $h(\bm{x})=\sum_{i=1}^{n}px_{i}-\frac{x_{i}^{p}}{1-p}$ ($10<p<$ 1);
  5. (v)
    Hellinger entropy kernel $h(\bm{x})=\sum_{i=1}^{n}-\sqrt{1-x_{i}^{2}}$.
Remark 1.
The separable structure of $h$ implies that $\operatorname{cl}(\operatorname{dom}(h))$ is a box of the form
$\operatorname{cl}(\operatorname{dom}(h))=[a,c]\times[a,c]\times\cdots\times[a,c],$
[…]
While Definition 1 imposes structural assumptions on the kernel function $h$, it also implies a continuity-like property that is frequently used in convergence analysis. Unlike many previous works that directly assume this as a technical condition, see, e.g., (De Pierro and Iusem, 1986, Definition 2.1 (vi)), (Chen and Teboulle, 1993, Definition 2.1 (v)), (Bauschke et al., 2017, Assumption H(ii)), (Byrne and Censor, 2001, B5, p. 95), (Souza et al., 2010, Definition 3.2 (B4)), and (Teboulle, 2018, Assumption H (iii)), we derive it here as a consequence of the separability and strict convexity properties of $h$.
Lemma 1.
Let $h:{\mathbb{R}}^{n}\rightarrow\overline{{\mathbb{R}}}$ be a separable kernel function. Suppose sequences $\{\bm{y}^{k}\}_{k\in{\mathbb{N}}},\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\rm int}(\operatorname{dom}(h))$ satisfy $\bm{z}^{k}\to\overline{\bm{z}}$ and $D_{h}(\bm{y}^{k},\bm{z}^{k})\to 0$. Then it follows that $\bm{y}^{k}\to\overline{\bm{z}}$.
Proof.
We prove the result by contradiction. Suppose, on the contrary, that $\bm{y}^{k}\not\to\overline{\bm{z}}$. Then, by passing to a subsequence if necessary, we may assume $y^{k}_{i_{0}}\to\overline{y}_{i_{0}}\in\operatorname{cl}(\operatorname{dom}(\varphi))$ with $\overline{y}_{i_{0}}\neq\overline{z}_{i_{0}}$ for some $i_{0}\in[n]$. Without loss of generality (WLOG), we may assume $\overline{z}_{i_{0}}<\overline{y}_{i_{0}}$. Then, there exist scalars $p,q\in{\rm int}(\operatorname{dom}(\varphi))$ such that $\overline{z}_{i_{0}}<p<q<\overline{y}_{i_{0}}$. By the convergence $z_{i_{0}}^{k}\to\overline{z}_{i_{0}}$ and $y_{i_{0}}^{k}\to\overline{y}_{i_{0}}$, there exists $k_{0}>0$ such that for all $k\geq k_{0}$, we have $z_{i_{0}}^{k}<p<q<y_{i_{0}}^{k}.$
By the three-point identity for Bregman divergences (see (Chen and Teboulle, 1993, Lemma 3.1)), for any $x,y,z\in\operatorname{dom}(\varphi)$ with $z\leq x\leq y$, we have
$D_{\varphi}(z,x)+D_{\varphi}(x,y)-D_{\varphi}(z,y)=(z-x)(\varphi^{\prime}(y)-\varphi^{\prime}(x))\leq 0,$
where the inequality follows from the strict convexity of $\varphi$. As a result, we obtain
$D_{\varphi}(z,x)\leq D_{\varphi}(z,y)\text{ and }D_{\varphi}(x,y)\leq D_{\varphi}(z,y),\quad\text{ for }z\leq x\leq y.$ (2)
It follows that $D_{h}(\bm{y}^{k},\bm{z}^{k})\geq D_{\varphi}(y^{k}_{i_{0}},z^{k}_{i_{0}})\geq D_{\varphi}(q,z^{k}_{i_{0}})\geq D_{\varphi}(q,p)>0$, which contradicts $D_{h}(\bm{y}^{k},\bm{z}^{k})\to 0$. We complete the proof. ∎
We now state the assumptions imposed on the optimization problem (P).
Assumption 1.
Let $\operatorname{dom}(g)={\cal X}$ be a nonempty closed convex set. We make the following assumptions:
  1. (i)
    The function $f$ is continuously differentiable on ${\cal X}$.
  2. (ii)
    The function $g$ is convex and locally Lipschitz continuous on ${\cal X}$.
  3. (iii)
    There exists a strictly feasible point $\bm{x}^{{\rm{int}}}\in{\rm{int}}(\operatorname{dom}(h))\cap{\cal X}$, and ${\cal X}\subseteq\operatorname{cl}(\operatorname{dom}(h))$.
  4. (iv)
    The function $h$ is a separable kernel function; see Definition 1.
Assumptions 1 (i)–(iii), or their stronger variants, are standard in the literature; see, e.g., (Bauschke et al., 2017, 2019, Assumption A), and (Azizian et al., 2022, Definition 1 and Assumption 1).
Next, we state the assumptions on the surrogate model $\gamma$ used in the update rule (1) of BPs.
Assumption 2 (Surrogate model $\gamma$).
The following conditions hold:
  1. (i)
    The mapping $(\bm{x},\bm{y})\mapsto\gamma(\bm{y};\bm{x})$, as well as the gradient mapping $(\bm{x},\bm{y})\mapsto\nabla\gamma(\bm{y};\bm{x})$ are jointly continuous with respect to $(\bm{y},\bm{x})$ for all $\bm{y}\in{\cal X}$ and $\bm{x}\in{\cal X}$.
  2. (ii)
    For all $\bm{x}\in{\cal X}$, we have $\nabla\gamma(\bm{y};\bm{x})\mid_{\bm{y}=\bm{x}}=\nabla f(\bm{x}),\text{ and }\gamma(\bm{y};\bm{x})\mid_{\bm{y}=\bm{x}}=f(\bm{x}).$
  3. (iii)
    There exists a constant $\overline{t}>0$ such that, for all $\bm{x}\in{\cal X}$, the function $\overline{t}(\gamma(\ \cdot\ ;\bm{x})+g(\cdot))+h(\cdot)$ is strictly convex.
  4. (iv)
    Either ${\cal X}$ is compact or the following condition holds: For all step sizes $t\in(0,\overline{t}]$ and all sequences $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}},\{\bm{y}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\rm int}(\operatorname{dom}(h))\cap{\cal X}$ with $\|\bm{y}^{k}\|\to+\infty$ and $\bm{z}^{k}\to\overline{\bm{z}}\in{\cal X}$, we have
    $\lim\limits_{k\to\infty}\gamma(\bm{y}^{k};\bm{z}^{k})+g(\bm{y}^{k})+\frac{1}{t}D_{h}(\bm{y}^{k},\bm{z}^{k})=+\infty.$ (3)
Unless otherwise specified, the step size $t$ in this paper is assumed to satisfy $t\in(0,\overline{t})$.
Assumptions 2 (i) and (ii) are standard, serving to ensure the continuity and local accuracy of the surrogate model $\gamma$. In all three choices of $\gamma$ discussed in the introduction, Assumption 2 (iii) is either a standard condition in the literature or is automatically satisfied. Specifically, when $\gamma$ is the first-order expansion of $f$ at the current iterate $\bm{x}$, Assumption 2 (iii) holds trivially. When $\gamma$ is taken as the original function $f$, the condition reduces to the relative convexity, a weaker assumption that has been extensively studied; see, e.g., (Bolte et al., 2018; Zhang and He, 2018). In the case where $\gamma$ is the second-order expansion of $f$, the $L$-smoothness of $f$ and the strict convexity of $h$ together suffice to guarantee Assumption 2 (iii). Assumption 2 (iv) ensures the well-posedness of the BPs. If ${\cal X}$ is compact, this condition holds automatically; see, e.g., (Bauschke et al., 2017, Lemma 2) and (Bolte et al., 2018, Assumption B). Interested readers are referred to Appendix A for the rigorous verification of this assumption for commonly used surrogate models.
Following (Teboulle, 2018, Lemma 2.3), we are now ready to state a key result concerning the well-posedness of the update rule (1) over ${\rm int}(\operatorname{dom}(h))$
Lemma 2.
Suppose that Assumptions 1 and 2 hold. Then for all $\bm{x}\in{\rm int}(\operatorname{dom}(h))\cap{\cal X}$, the update mapping $T_{\gamma}^{t}:{\mathbb{R}}^{d}\rightarrow{\mathbb{R}}^{d}$ defined by
$T_{\gamma}^{t}(\bm{x}):=\operatorname*{argmin}_{\bm{y}\in{\cal X}}\left\{\gamma(\bm{y};\bm{x})+g(\bm{y})+\frac{1}{t}D_{h}(\bm{y},\bm{x})\right\}$
is well-defined, and satisfies $T_{\gamma}^{t}(\bm{x})\in{\rm int}(\operatorname{dom}(h))\cap{\cal X}.$
Proof.
By Assumption 2 (iv), for any sequence $\{\bm{y}^{k}\}\subset{\rm int}(\operatorname{dom}(h))\cap\mathcal{X}$ with $\|\bm{y}^{k}\|\to\infty$ and $\bm{x}\in{\cal X}$, it holds that
$\lim_{k\to\infty}\left\{\gamma(\bm{y}^{k};\bm{x})+g(\bm{y}^{k})+\frac{1}{t}D_{h}(\bm{y}^{k},\bm{x})\right\}=+\infty.$
This implies that the objective function is coercive over ${\rm int}(\operatorname{dom}(h))\cap\mathcal{X}$, i.e., it tends to infinity as $\|\bm{y}\|\to\infty$. Hence, there exists a radius $r>0$ such that the infimum is attained over the compact set $\mathcal{X}\cap{\mathbb{B}}_{r}({\mathbf{0}})$. That is,
$\inf_{\bm{y}\in\mathcal{X}}\left\{\gamma(\bm{y};\bm{x})+g(\bm{y})+\frac{1}{t}D_{h}(\bm{y},\bm{x})\right\}=\inf_{\bm{y}\in\mathcal{X}\cap{\mathbb{B}}_{r}({\mathbf{0}})}\left\{\gamma(\bm{y};\bm{x})+g(\bm{y})+\frac{1}{t}D_{h}(\bm{y},\bm{x})\right\}>-\infty.$
Since the objective function is continuous over ${\cal X}$ and the feasible set is closed and bounded, the infimum is attained. By the definition of $T_{\gamma}^{t}$ and Teboulle (2018, Lemma 2.3 (a)), we conclude that $T_{\gamma}^{t}(\bm{x})$ is well-defined and lies in $\operatorname{int}(\operatorname{dom}(h))\cap\mathcal{X}$. ∎
At last, to evaluate how close a given iterate is to stationarity, it is common to introduce a residual mapping $R:\mathbb{R}^{n}\to\mathbb{R}_{+}$ that quantifies the degree of stationarity. Such a residual mapping plays a central role not only in convergence analysis but also in the design of practical stopping criteria. In this paper, we adopt the following Bregman divergence-based stationarity measure:
$R_{\gamma}^{t}(\bm{x}):=D_{h}\left(T^{t}_{\gamma}(\bm{x}),\bm{x}\right),$ (4)
which measures the discrepancy between the current iterate $\bm{x}$ and its update $T^{t}_{\gamma}(\bm{x})$ under the Bregman geometry induced by $h$. This formulation unifies various residual-type stationarity measures commonly used in the analysis of BPs; see, e.g., (Bedi et al., 2022; Huang et al., 2022a, b; Latafat et al., 2022). In particular, if we set $\gamma=f$, then the update mapping $T^{t}_{\gamma}(\bm{x})$ reduces to the standard Bregman proximal operator, and $R_{\gamma}^{t}(\bm{x})$ coincides with the stationarity gap $D_{h}(\operatorname{prox}_{h,F}^{t}(\bm{x}),\bm{x})$ introduced by Zhang and He (2018). To make the connection precise, we recall the definition of the Bregman proximal mapping:
Definition 2 (Bregman proximal mapping (Bauschke et al., 2018; Lau and Liu, 2022)).
Let $F:{\rm int}(\operatorname{dom}(h))\to{\mathbb{R}}$ and $t>0$. The Bregman proximal mapping for $F$ associated with the kernel $h$ is defined by
${\operatorname{prox}}^{t}_{h,F}(\bm{x})=\underset{\bm{y}\in{\mathbb{R}}^{n}}{\operatorname*{argmin}}\left\{F(\bm{y})+\frac{1}{t}D_{h}(\bm{y},\bm{x})\right\}.$
Note that the function $\gamma$ in (4) is independent of the algorithmic update and can be chosen differently. For example, as in Zhang and He (2018), in the analysis of BPG, one may let $\gamma=f$ in (4) even if the algorithm uses a linear approximation of $f$ as $\gamma$ instead.
3 Main Results
In this section, we present our three main theoretical results. We begin by providing a complete characterization of the stationarity measure introduced in (4). To state our main results precisely, we first define the following index sets associated with a point $\bm{x}\in{\mathbb{R}}^{n}$:
${\cal B}(\bm{x}):=\left\{b\in[n]:x_{b}\in{\rm{bd}}\left(\operatorname{dom}(\varphi)\right)\right\},\quad{\cal I}(\bm{x}):=\left\{i\in[n]:x_{i}\in{\rm{int}}\left(\operatorname{dom}(\varphi)\right)\right\}.$
Here, ${\cal B}(\bm{x})$ and ${\cal I}(\bm{x})$ denote the sets of coordinate indices where $\bm{x}$ lies on the boundary and in the interior of $\operatorname{dom}(\varphi)$, respectively.
Theorem 1.
Let $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\rm int}(\operatorname{dom}(h))\cap{\cal X}$ be a sequence that converges to $\overline{\bm{z}}\in{\cal X}$. Then the following equivalence holds:
$\lim\limits_{k\to\infty}R^{t}_{\gamma}(\bm{z}^{k})=0\ \Longleftrightarrow\ \exists\ \bm{p}\in\partial F(\overline{\bm{z}})\text{ such that }\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}.$
Theorem 1 shows that, for any feasible sequence, convergence of the stationarity measure to zero does not guarantee that the limit point is stationary. Instead, it only ensures the existence of a subgradient at the limit point whose interior coordinates vanish. As a result, one cannot conclude that the output of a BP algorithm is approximately stationary—even when the residual $R_{\gamma}^{t}$ is small.
Importantly, this limitation is not a consequence of the stationarity measure itself, but rather reflects a more fundamental issue. The residual measure $R_{\gamma}^{t}$ captures the descent behavior inherent to BPs, and has been widely investigated in nonconvex settings; see, e.g., (Bauschke et al., 2017; Zhang and He, 2018). In such settings, BPs are typically only guaranteed to achieve sufficient descent with respect to the Bregman divergence. Beyond this, little can be rigorously said about the behavior of their iterates. As a result, Theorem 1 implies that no further guarantees, such as approximate stationarity, can be obtained under current algorithmic frameworks based on Bregman divergence.
We formalize this conceptual limitation by introducing the notion of spurious stationary points, which naturally emerge from the equivalence established in Theorem 1.
Definition 3 (Spurious stationary points).
A point $\bm{x}\in{\cal X}$ is defined as a spurious stationary point of problem (P) if there exists a vector $\bm{p}\in\partial F(\bm{x})$ satisfying $\bm{p}_{{\cal I}(\bm{x})}=0$ but ${\mathbf{0}}\notin\partial F(\bm{x})$.
Remark 2.
Spurious stationary points can only arise when the kernel $h$ is not gradient Lipschitz continuous. Indeed, if $h$ has Lipschitz continuous gradient, then by Definition 1 (ii), we have $\operatorname{dom}(h)={\mathbb{R}}^{n}$ and ${\cal I}(\bm{x})=[n]$ hold for all $\bm{x}\in{\cal X}$, which rules out the possibility of spurious stationary points by definition.
While the previous discussion exposes the conceptual limitation of the Bregman divergence-based stationarity measure (4), the following result shows that this limitation is algorithmically unavoidable. Specifically, once a spurious stationary point exists, BPs can become trapped arbitrarily close to it for any finite number of iterations, regardless of the choice of stationarity measure. This phenomenon reveals that the difficulty lies not in how stationarity is quantified, but in the structural behavior of the algorithm itself.
Theorem 2 (Hardness).
Suppose that there exists a spurious stationary point $\tilde{\bm{x}}^{\star}\in{\cal X}$ for problem (P). For every $K\in{\mathbb{N}}$ and $\epsilon>0$, there exists an initial point $\bm{x}^{0}\in{\mathbb{B}}_{\epsilon}(\tilde{\bm{x}}^{\star})\cap{\cal X}\cap{\rm int}(\operatorname{dom}(h))$, sufficiently close to the spurious stationary point $\tilde{\bm{x}}^{\star}$, such that the sequence $\{\bm{x}^{k}\}_{k\in[K]}$ generated by (1) satisfies
$\bm{x}^{k}\in{\mathbb{B}}_{\epsilon}(\tilde{\bm{x}}^{\star})\quad\text{ for all }k\in[K].$ (5)
Proof.
By Theorem 1 and Definition 3, for every sequence $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\rm int}(\operatorname{dom}(h))\cap{\cal X}$ converging to a spurious stationary point $\tilde{\bm{x}}^{\star}$, we have $\lim_{k\to\infty}R_{\gamma}^{t}(\bm{z}^{k})=\lim_{k\to\infty}D_{h}(T^{t}_{\gamma}(\bm{z}^{k}),\bm{z}^{k})=0$. Moreover, under Assumption 1 (iv), Lemma 1 implies that $\lim_{k\to\infty}{T}^{t}_{\gamma}(\bm{z}^{k})=\tilde{\bm{x}}^{\star}$. Since both sequences $\{\bm{z}^{k}\}_{k\geq 0}$ and $\{T^{t}_{\gamma}(\bm{z}^{k})\}_{k\geq 0}$ converges to $\tilde{\bm{x}}^{\star}$, it follows that for any $\epsilon>0$, there exists $\delta>0$ such that for all $\bm{x}\in{\cal X}\cap{\rm int}(\operatorname{dom}(h))$ with $\|\bm{x}-\tilde{\bm{x}}^{\star}\|<\delta$, we have $\|T^{t}_{\gamma}(\bm{x})-\tilde{\bm{x}}^{*}\|<\epsilon$. Then, we can let $\bm{x}=\bm{x}^{K-1}$ and choose a small constant $\epsilon_{1}<\min\{\delta,\frac{1}{2}\epsilon_{0}\}$ such that
$\|\bm{x}^{K}-\tilde{\bm{x}}^{\star}\|=\|T^{t}_{\gamma}(\bm{x}^{K-1})-\tilde{\bm{x}}^{\star}\|\leq\epsilon_{0},$
whenever $\|\bm{x}^{K-1}-\tilde{\bm{x}}^{\star}\|<\epsilon_{1}$.
Repeating the above argument for $K$ iterations, we inductively construct a sequence $\{\epsilon_{k}\}_{k=0}^{K}$ such that $\epsilon_{k+1}\leq\tfrac{1}{2}\epsilon_{k}$. Then, for any $k=0,1,\dots,K$, we have $\|\bm{x}^{K-k}-\tilde{\bm{x}}^{\star}\|\leq\epsilon_{k}$ provided that $\|\bm{x}^{K-k-1}-\tilde{\bm{x}}^{\star}\|\leq\epsilon_{k+1}.$ That is, if the initial point $\bm{x}^{0}=\bm{x}^{K-K}$ satisfies $\|\bm{x}^{0}-\tilde{\bm{x}}^{\star}\|\leq\epsilon_{K}$, then all subsequent iterates up to $\bm{x}^{K}$ remain within ${\mathbb{B}}_{\epsilon_{0}}(\tilde{\bm{x}}^{\star})$. We finished the proof. ∎
While Theorem 2 establishes the algorithmic hardness caused by spurious stationary points, it naturally raises a critical question: Are such points common in practice, or are they merely pathological artifacts? The next result shows that spurious stationary points are not merely pathological artifacts. In fact, they are ubiquitous in a broad class of nonconvex problems with polyhedral constraints.
Theorem 3.
Consider the optimization problem
$\min_{\bm{x}\in\mathcal{X}}~{}f(\bm{x}),\quad\text{where }\mathcal{X}:=\{\bm{x}\in{\mathbb{R}}^{n}:\mathbf{A}\bm{x}=\mathbf{b},\ \bm{x}\geq{\mathbf{0}}\}$
is not singleton, with $\mathbf{A}\in{\mathbb{R}}^{m\times n}$ and $\mathbf{b}\in\mathbb{R}^{m}$. Assume that $\operatorname{dom}(\partial f)\supseteq{\cal X}$, and that the kernel function $h$ satisfies $\operatorname{cl}(\operatorname{dom}(h))={\mathbb{R}}^{n}_{+}$. Then, every non-stationary vertex of $\mathcal{X}$ is a spurious stationary point.
Proof.
At any vertex $\bm{x}$ of ${\cal X}$, the active constraint gradients consist of the rows of $\mathbf{A}$ together with the standard basis vectors $\bm{e}_{i}$ for all $i\in{\cal B}(\bm{x})$. These vectors together span $\mathbb{R}^{n}$, that is,
$\operatorname{rank}\left(\left[\mathbf{A}^{T},\bm{e}_{i}:i\in{\cal B}(\bm{x})\right]\right)=n,$
as shown in (Schrijver, 1998, Section 8.5). Since ${\cal X}$ is not a singleton, we must have $\operatorname{rank}(\mathbf{A})<n$. Hence, any vertex $\bm{x}\in{\cal X}$ must have at least one active inequality constraint, i.e., $\mathcal{B}(\bm{x})\neq\emptyset$.
Now, fix any $\bm{v}\in\partial f(\bm{x})$. By the above rank condition and the assumption $\operatorname{dom}(\partial f)\supseteq{\cal X}$, there exist vectors $\bm{\mu}\in\mathbb{R}^{m}$ and $\bm{\lambda}\in\mathbb{R}^{|{\cal B}(\bm{x})|}$ such that
$\bm{v}+\mathbf{A}^{T}\bm{\mu}+\sum_{i\in{\cal B}(\bm{x})}\lambda_{i}\bm{e}_{i}={\mathbf{0}},$
which implies that $(\bm{v}+\mathbf{A}^{T}\bm{\mu})_{{\cal I}(\bm{x})}={\mathbf{0}}.$ Let $F(\bm{x}):=f(\bm{x})+\delta_{{\cal X}}(\bm{x})$. If the vertex $\bm{x}$ is not a stationary point of $F$, then by definition there exists a vector
$\bm{p}:=\bm{v}+\mathbf{A}^{\top}\bm{\mu}\in\partial F(\bm{x})\quad\text{such that}\quad\bm{p}_{{\cal I}(\bm{x})}=\mathbf{0},\quad\text{but}\quad\mathbf{0}\notin\partial F(\bm{x}).$
Thus, by Definition 3, $\bm{x}$ is a spurious stationary point.
4 Practical Implication of Hardness Results
Theorem 2 reveals a fundamental limitation of BPs: Even when the stationarity measure vanishes, the iterates can remain arbitrarily close to a spurious stationary point for any finite number of steps.
This section explores the practical implications of this phenomenon through two illustrative examples. The first revisits the LP instance (c-ex) and shows that Theorem 2 does not contradict known non-asymptotic convergence guarantees. However, it highlights that convergence speed can depend critically on initialization. The second example focuses on a nonconvex objective and illustrates that, unlike in convex settings, initializing well within the interior of the kernel’s domain does not guarantee avoidance of pathological behaviors. As iterations progress, BPs may still drift toward a spurious stationary point and become trapped. Together, these examples underscore the subtle yet widespread risk posed by spurious stationary points under Bregman geometry.
We first examine the LP instance given in (c-ex).
Example 2.
Suppose that $\operatorname{cl}(\operatorname{dom}(h))={\mathbb{R}}^{2}_{+}$ and consider the problem
$\begin{array}[]{rl}\min&-x_{1}\\ \text{s.t.}&x_{1}+x_{2}=1,\quad x_{1},x_{2}\geq 0.\end{array}$
A simple calculation reveals that the point $\tilde{\bm{x}}^{\star}=(0,1)$ is not a stationary point:
${\mathbf{0}}\notin\partial F(\tilde{\bm{x}}^{\star})$ =$=\left\{(-1,0)+\lambda(-1,0)+\mu(1,1):\lambda\in{\mathbb{R}}_{+},\mu\in\mathbb{R}\right\}.$
Moreover, the interior coordinate set at $\tilde{\bm{x}}^{\star}$ is $\mathcal{I}(\tilde{\bm{x}}^{\star})=\{2\}$, and there exists a subgradient $\bm{p}\in\partial F(\tilde{\bm{x}}^{\star})$ such that $\bm{p}_{\mathcal{I}(\tilde{\bm{x}}^{\star})}=0$. By Definition 3, we conclude that $\tilde{\bm{x}}^{\star}$ is a spurious stationary point.
To illustrate Theorem 2, we adopt the Boltzmann–Shannon entropy kernel $\varphi(x)=x\log x$, which is widely used in constrained optimization problems over the simplex. Under this kernel, BPG update reduces to the classical multiplicative weights update method (Arora et al., 2012):
$\bm{x}^{k+1}$ $=\operatorname*{argmin}_{\bm{y}\in{\mathbb{R}}^{2}}t(-1,0)^{T}\bm{y}+D_{h}(\bm{y},\bm{x}^{k})+\delta_{\Delta_{2}}(\bm{y})$
$=\left(\frac{x_{1}^{k}}{x_{1}^{k}+e^{-t}x_{2}^{k}},\frac{e^{-t}x_{2}^{k}}{x_{1}^{k}+e^{-t}x_{2}^{k}}\right),\quad\forall k\in[K].$
We initialize the algorithm with
$\bm{x}^{0}=\left(\frac{\sqrt{2}\epsilon}{2}e^{-tK},1-\frac{\sqrt{2}\epsilon}{2}e^{-tK}\right),$
which lies strictly inside the simplex. Then, it is straightforward to verify that for all $k\in[K]$,
$\|\bm{x}^{k+1}-\tilde{\bm{x}}^{\star}\|=\frac{\sqrt{2}x_{1}^{k}}{x_{1}^{k}+e^{-t}x_{2}^{k}}\leq\sqrt{2}e^{t}x_{1}^{k}\leq\sqrt{2}e^{tk}x_{1}^{0}=e^{-t(K-k)}\epsilon\leq\epsilon,$
where the first inequality is derived from the constraint $x_{1}^{k}+x_{2}^{k}=1$ and $t\geq 0$, the second inequality is justified by iteratively applying the recursive relation from the first inequality $k$ times.
This example shows that when the initial point is extremely close to a spurious stationary point, the trapping phenomenon described in Theorem 2 can be triggered. Importantly, this behavior does not contradict the non-asymptotic convergence guarantee established in Bauschke et al. (2017, Corollary 1). That result states that the sequence $\{\bm{x}^{k}\}_{k\in\mathbb{N}}$ generated by BPG satisfies:
$f(\bm{x}^{K})-\min_{\bm{x}\in{\cal X}}\ f\leq\frac{D_{h}(\bm{x}^{\star},\bm{x}^{0})}{t}\cdot\frac{1}{K},$ (6)
where $\bm{x}^{\star}\in\operatorname*{argmin}_{\bm{x}\in{\cal X}}f$ is the global minimizer, $t$ is the step size, and $\bm{x}^{0}$ is the initial point. Here, the rate depends linearly on the initial Bregman divergence $D_{h}(\bm{x}^{\star},\bm{x}^{0})$, which can be made arbitrarily large by placing $\bm{x}^{0}$ near a spurious stationary point. For example, in the LP instance above, choosing $\bm{x}^{0}=(\exp(-K),1-\exp(-K))$ yields $D_{h}((1,0),(\exp(-K),1-\exp(-K)))=K$, leading to the trivial upper bound:
$f(\bm{x}^{K})-\min_{\bm{x}\in{\cal X}}f(\bm{x})\leq\frac{1}{t}.$
This illustrate that, while the objective gap converges asymptotically, the iterates may stay in suboptimal regions — far away from the global optimal — for an arbitrarily long time.
The above LP instance may give the impression that pathological behavior only arises when the initialization is extremely close to a spurious stationary point, which, in general, lies precisely on the boundary of $\operatorname{dom}(h)$. However, this intuition is largely valid only for convex problems. In contrast, for nonconvex problems, this intuition fails: Even when the algorithm is initialized well within the interior of the domain and far away from any spurious stationary points, the iterates may gradually drift toward a spurious stationary point. Once sufficiently close, the conditions of Theorem 2 may be activated dynamically during the optimization process, leading to stagnation. The following example illustrates precisely this behavior: Despite a seemingly benign initialization, the iterates eventually become trapped near a spurious stationary point.
Example 3.
Suppose that $\operatorname{cl}(\operatorname{dom}(h))={\mathbb{R}}^{2}_{+}$ and consider the following parameterized nonconvex optimization problem:
$\begin{array}[]{rl}\min\limits_{\bm{x}\in{\mathbb{R}}^{2}}&\ f_{\alpha}(x_{1},x_{2}):=\frac{1}{2}\phi_{\alpha}(x_{1})\cdot(x_{2}+0.05)\\ {\rm s.t.}&x_{1},x_{2}\in[0,1],\end{array}$ (7)
where $\alpha\in(0,0.1]$ and $\phi:[0,1]\rightarrow{\mathbb{R}}$ is a continuously differentiable function satisfying
  1. (i)
    $\phi_{\alpha}(x)=2(x-\alpha)$ for $x\geq a$, $\phi_{\alpha}(x)\leq 0$ for $x\leq\alpha$, and $\phi_{\alpha}(0)\leq-1$;
  2. (ii)
    $\phi^{\prime}_{\alpha}(x)\geq 1$ for $x\in[0,1]$;
  3. (iii)
    $f_{\alpha}+h$ and $-f_{\alpha}+h$ are convex on $(0,1]\times(0,1]$.
First of all, we compute the first-order optimality condition of (7) as
$\partial F(\bm{x})=\left\{\left(\frac{1}{2}\phi^{\prime}_{\alpha}(x_{1})(x_{2}+0.05),\frac{1}{2}\phi_{\alpha}(x_{1})\right)+\bm{\lambda}-\bm{\mu}:\bm{\lambda},\bm{\mu}\geq{\mathbf{0}},\bm{\lambda}^{\top}({\mathbf{1}}-\bm{x})=0,\bm{\mu}^{\top}\bm{x}=0\right\}.$ (8)
Combining the subdifferential characterization in (8) with condition (ii) in Example 3, we observe that for any $\bm{x}\in(0,1]\times[0,1]$, all subgradients $\bm{p}\in\partial F(\bm{x})$ satisfy $p_{1}\geq 0.025$. This implies that any true stationary or spurious stationary point must lie on the boundary $x_{1}=0$. Substituting $x_{1}=0$ into (8) and invoking Definition 3, we find that $\tilde{\bm{x}}^{\star}=(0,0)$ is a spurious stationary point, while $\bm{x}^{\star}=(0,1)$ is the unique true stationary point and hence the global minimizer.
Importantly, from the above discussion, it follows directly that $\operatorname{dist}(\bm{0},\partial F(\bm{x}))\geq 0.025$ for all $\bm{x}\in[0,1]\times[0,1]\setminus\{\bm{x}^{\star}\}$. This sharpness property (Burke and Ferris, 1993) implies that the vanilla Projected Gradient Descent (PGD) method achieves a uniform decrease in the objective value at every iterate except the global minimizer, and may converge in a finite number of steps. Moreover, condition (iii) in Example 3 is a commonly adopted assumption to guarantee non-asymptotic convergence rates for BPG under the stationarity measure (4); see, e.g., (Zhang and He, 2018, Theorem 4.1) and (Bolte et al., 2018, Proposition 4.1).
Given these favorable properties, the problem in Example 3 may initially appear benign: It admits a unique global minimizer, enjoys a sharpness property that favors fast convergence under PGD, and satisfies structural conditions often used to guarantee non-asymptotic convergence for BPG with the measure (4). It is thus natural to expect that initializing sufficiently far from the spurious stationary point should guarantee non-asymptotic convergence to the global optimum.
However, this intuition proves to be misleading. Despite the seemingly favorable properties discussed above, the iterates generated by BPG can still be drawn toward the spurious stationary point $\tilde{\bm{x}}^{\star}=(0,0)$ and become trapped. We begin by outlining the high-level intuition behind this phenomenon, followed by a concrete numerical example that empirically illustrates the pathological behavior.
We now instantiate the BPG method using the Burg entropy kernel $\varphi(x)=-\log(x)$, and derive its explicit update rule by applying the first-order optimality condition to (7):
[…]
Since we observe that $\nabla_{x_{1}}f_{\alpha}(x_{1}^{k},x_{2}^{k})=\frac{1}{2}\phi_{\alpha}^{\prime}(x_{1}^{k})(x_{2}^{k}+0.05)>0$, the BPG update in the $x_{1}$-coordinate yields $x_{1}^{k+1}<x_{1}^{k}\leq 1$. This implies that the iterates are monotonically decreasing in $x_{1}$ and will continue contracting toward zero. The behavior along the $x_{2}$-coordinate, however, is more subtle. The first-order optimality condition leads to the explicit update:
[…]
When $\alpha$ is small, during the early iterations with $x_{1}^{k}>\alpha$, condition (i) in Example 3 guarantees that $\phi_{\alpha}(x_{1}^{k})>0$. As a result, both $x_{1}^{k}$ and $x_{2}^{k}$ decrease simultaneously, driving the iterates toward the origin. However, once $x_{1}^{k}<\alpha$, the function $\phi_{\alpha}$ becomes nonpositive, and the $x_{2}$-update begins to reverse direction. Unfortunately, by this point, the iterates may have already been drawn sufficiently close to the spurious stationary point $(0,0)$, making it too late for the algorithm to recover. As a result, the sequence becomes trapped in a neighborhood of the spurious point.
We are now ready to present an explicit construction of the function $\phi_{\alpha}$ that satisfies the conditions in Example 3. Specifically, we define $\phi_{\alpha}:[0,1]\rightarrow\mathbb{R}$ as the following piecewise continuously differentiable function:
$\phi_{\alpha}(x)=\begin{cases}2x-2\alpha,&\text{if }x\in[\alpha,1];\\ x+\alpha\log\left(\frac{x}{\alpha}\right)-\alpha,&\text{if }x\in[\alpha\exp(-\frac{1}{\alpha}),\alpha];\\ x-\alpha\log\left(\frac{x}{\alpha}\right)+2\exp(\frac{1}{\alpha})x-3\alpha-2,&\text{if }x\in[\frac{1}{2}\alpha\exp(-\frac{1}{\alpha}),a\exp(-\frac{1}{\alpha})];\\ x+\alpha(\log 2-2)-1,&\text{if }x\in[0,\frac{1}{2}\alpha\exp(-\frac{1}{\alpha})].\end{cases}$ (11)
To visualize the pathological behavior in Example 3, we set $\alpha=0.01$, use a maximum iteration count of 5000 and initialize the iterates at $(x_{1}^{0},x_{2}^{0})=(1,0.1)$, which is far from both the boundary of the domain and the spurious stationary point $(0,0)$. Figure 2 compares the dynamics of PGD and BPG on this instance. In Figure 2 (c), we observe that PGD rapidly converges to the global minimum $(0,1)$ in fewer than 50 iterations. In stark contrast, the BPG trajectory is gradually drawn toward the spurious stationary point $(0,0)$, where it becomes trapped for more than 5000 iterations. Figure 2 (a) and (b) illustrate the gradient fields under Euclidean and Bregman geometries, respectively. Notably, in Figure 2 (b), we see that within the region $x_{1}>\alpha$, BPG update directions consistently point toward $(0,0)$, making the spurious stationary point an attractor. Once the iterate enters the zone $x_{1}<\alpha$, the dynamics flip abruptly due to the sign change in $\phi_{\alpha}(x_{1})$, but by then the algorithm has already entered the trap. This attract-and-flip behavior arises fundamentally from the nonconvexity of the problem and cannot be avoided under the Bregman geometry.
At last, we summarize the above behavior in the following formal observation, which demonstrates that the BPG iterates can remain arbitrarily close to a spurious stationary point for any prescribed number of iterations.
Observation 1.
Consider Example 3, where the objective function $\phi_{\alpha}$ is specified as in (11). We apply BPG with the Burg entropy kernel $\varphi(x)=-\log(x)$, using a constant stepsize $t=0.5$ and initialization $\bm{x}^{0}=(1,0.1)$. Let $\{\bm{x}^{k}\}_{k\in{\mathbb{N}}}$ denote the resulting sequence of iterates. Then, for any given $K\in\mathbb{N}$ and $\epsilon\in(0,1]$, there exists a parameter $\alpha>0$ and an iteration index $k_{1}>0$ such that the iterates satisfy
$\bm{x}^{k}\in\mathbb{B}_{\epsilon}(\tilde{\bm{x}}^{\star}),\quad\text{for all }k_{1}\leq k\leq K+k_{1}.$
The rigorous proof of this result is provided in Appendix B.
Figure 2: Comparison of PGD and BPG dynamics with the Burg entropy kernel on the nonconvex instance constructed in Example 3, where $\phi_{\alpha}$ is given in (11). (a) Euclidean gradient field $-\nabla f_{\alpha}(\bm{x})$. (b) Bregman update field induced by Burg entropy, i.e., $T_{\gamma}^{1}(\bm{x})-\bm{x}$. (c) Trajectories of BPG and PGD starting from $(1,0.1)$. While PGD quickly converges, BPG becomes trapped near the spurious stationary point $(0,0)$.
5 Proof of Theorem 1
We now present the proof of Theorem 1. The overall structure follows the diagram below:
$\exists\ \bm{p}\in\partial F(\overline{\bm{z}})\text{ s.t. }\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}$$\overline{R}^{t}_{\gamma}(\overline{\bm{z}})=0$$\lim\limits_{k\to\infty}{R}^{t}_{\gamma}(\bm{z}^{k})=0$Prop. 1 Prop. 2
The full argument proceeds in three main steps:
  • (1)
    We define the extended Bregman stationarity measure $\overline{R}_{\gamma}^{t}$, which is well-defined on the entire domain ${\cal X}$, and satisfies $\overline{R}^{t}_{\gamma}(\bm{x})={R}^{t}_{\gamma}(\bm{x})$ for $\bm{x}\in{\cal X}\cap{\rm int}(\operatorname{dom}(h))$. This measure serves as a technical bridge between the limiting behavior of the original stationarity measure $R_{\gamma}^{t}(\bm{z}^{k})$ and the variational condition that there exists $\bm{p}\in\partial F(\overline{\bm{z}})$ satisfying $\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}$.
  • (2)
    We establish that $\overline{R}_{\gamma}^{t}(\overline{\bm{z}})=0$ if and only if there exists $\bm{p}\in\partial F(\overline{\bm{z}})$ with $\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}$.
    Proposition 1.
    For all $\overline{\bm{z}}\in{\cal X}$, the extended stationarity measure equals zero, i.e., $\overline{R}^{t}_{\gamma}(\overline{\bm{z}})=0$, if and only if there exists a vector $\bm{p}\in\partial F(\overline{\bm{z}})$ such that $\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}$.
    The proof of Proposition 1 are referred to Sec. 5.2.
  • (3)
    Finally, we prove that the extended Bregman stationarity measure is continuous.
    Proposition 2.
    The extended stationarity measure $\overline{R}^{t}_{\gamma}:\mathcal{X}\rightarrow\mathbb{R}$ is continuous on the domain $\mathcal{X}$.
    We defer the proof of Proposition 2 to Section 5.3.
By Proposition 2 and $\overline{R}^{t}_{\gamma}(\bm{x})={R}^{t}_{\gamma}(\bm{x})$ for $\bm{x}\in{\cal X}\cap{\rm int}(\operatorname{dom}(h))$, we see that for any sequence $\{\bm{z}^{k}\}_{k\geq 0}\subseteq{\cal X}\cap{\rm int}(\operatorname{dom}(h))$ converging to $\overline{\bm{z}}\in{\cal X}$,
$\lim_{k\to\infty}R^{t}_{\gamma}(\bm{z}^{k})=\lim_{k\to\infty}\overline{R}^{t}_{\gamma}(\bm{z}^{k})=\overline{R}^{t}_{\gamma}(\overline{\bm{z}}).$
Combined with Proposition 1, we complete the proof of Theorem 1.
We now present detailed arguments for each of the above components.
5.1 Extended Bregman Stationarity Measure
To address the issue that the original stationarity measure $R_{\gamma}^{t}$ is undefined on the boundary of $\operatorname{dom}(h)$, we first define an extended Bregman update mapping over the entire domain ${\cal X}$. This yields an extended Bregman stationarity measure, denoted by $\overline{R}_{\gamma}^{t}$, which serves as a central object in our subsequent analysis.
Definition 4 (Extended Bregman Update Mapping).
We define the extended Bregman update mapping $\overline{T}_{\gamma}^{t}(\bm{x})$ for all $\bm{x}\in\mathcal{X}$ by
$\overline{T}_{\gamma}^{t}(\bm{x}):=\operatorname*{argmin}_{\bm{y}\in\mathcal{X}}\;G_{\gamma}^{t}(\bm{y};\bm{x}),$
where the objective $G_{\gamma}^{t}(\bm{y};\bm{x})$ is given by
$G^{t}_{\gamma}(\bm{y};\bm{x}):=\gamma(\bm{y};\bm{x})+g(\bm{y})+\underbrace{\frac{1}{t}\sum_{i\in{\cal I}(\bm{x})}D_{\varphi}(y_{i},x_{i})}_{{\rm Interior\ coordinates}}+\underbrace{\delta_{\bm{y}_{{\cal B}(\bm{x})}=\bm{x}_{{\cal B}(\bm{x})}}(\bm{y})}_{{\rm Boundary\ coordinates}}.$
Here, the indicator term enforces the coordinates corresponding to the boundary indices ${\cal B}(\bm{x})$ remain fixed, while BPs is applied only over the interior coordinates ${\cal I}(\bm{x})$.
To establish the well-posedness of the extended Bregman update mapping $\overline{T}^{t}_{\gamma}$, we first verify that the objective function $G^{t}_{\gamma}(\bm{y};\bm{x})$ is coercive in $\bm{y}$. The following lemma formalizes this property.
Lemma 3 (Coerciveness of $G$).
Suppose that the sequence $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\cal X}$ converges to $\overline{\bm{z}}\in{\cal X}$ and the sequence $\{\bm{y}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\cal X}$ satisfies $\|\bm{y}^{k}\|\to+\infty$. Then, we have
$\lim_{k\to\infty}G^{t}_{\gamma}(\bm{y}^{k};\bm{z}^{k})=+\infty.$
Proof.
WLOG, we may assume that the boundary coordinates of $\bm{y}^{k}$ and $\bm{z}^{k}$ agree, i.e., $\bm{y}^{k}_{{\cal B}(\bm{z}^{k})}\equiv\bm{z}^{k}_{{\cal B}(\bm{z}^{k})}$; Otherwise, the indicator term in $G_{\gamma}^{t}(\bm{y}^{k};\bm{z}^{k})$ equals +$\infty$+ ∞ for some $k\in{\mathbb{N}}$, and the result trivially holds.
Let $\bm{z}^{{\rm int}}\in{\rm int}(\operatorname{dom}(h))\cap{\cal X}$ be any interior point. For each $k\in{\mathbb{N}}$, we define
$\hat{\bm{z}}^{k}=(1-\theta_{k})\bm{z}^{k}+\theta_{k}\bm{z}^{{\rm int}},\quad\hat{\bm{y}}^{k}=(1-\theta_{k})\bm{y}^{k}+\theta_{k}\bm{z}^{{\rm int}},$
where $\theta_{k}\in(0,1)$ with $\theta_{k}\to 0$. Then, it is easy to verify that $\hat{\bm{z}}^{k}\to\overline{\bm{z}}$, $\|\hat{\bm{y}}^{k}\|\to+\infty$, and $\hat{\bm{y}}^{k}_{{\cal B}(\bm{z}^{k})}=\hat{\bm{z}}^{k}_{{\cal B}(\bm{z}^{k})}$ for all $k\in{\mathbb{N}}$. It implies that
$\gamma(\hat{\bm{y}}^{k};\hat{\bm{z}}^{k})+g(\hat{\bm{y}}^{k})+\frac{1}{t}D_{h}(\hat{\bm{y}}^{k},\hat{\bm{z}}^{k})=\gamma(\hat{\bm{y}}^{k};\hat{\bm{z}}^{k})+g(\hat{\bm{y}}^{k})+\frac{1}{t}\sum_{i\in{\cal I}(\bm{z}^{k})}D_{\varphi}(\hat{y}^{k}_{i},\hat{z}^{k}_{i}).$
By the continuity of $\gamma,g,D_{\varphi}$, and the fact that $\theta_{k}\to 0$, we can choose $\theta_{k}$ small enough so that
$\left|\gamma(\hat{\bm{y}}^{k};\hat{\bm{z}}^{k})+g(\hat{\bm{y}}^{k})+\frac{1}{t}D_{h}(\hat{\bm{y}}^{k},\hat{\bm{z}}^{k})-G^{t}_{\gamma}(\bm{y}^{k};\bm{z}^{k})\right|\leq 1.$
Now, by Assumption 2 (iv), we know $\gamma(\hat{\bm{y}}^{k};\hat{\bm{z}}^{k})+g(\hat{\bm{y}}^{k})+\frac{1}{t}D_{h}(\hat{\bm{y}}^{k},\hat{\bm{z}}^{k})\to+\infty$. Consequently, we can conclude that $G^{t}_{\gamma}(\bm{y}^{k};\bm{z}^{k})\to+\infty$. We finished the proof. ∎
With Lemma 3 in place, we now present key properties of the extended Bregman update mapping $\overline{T}_{\gamma}^{t}$. These properties are central to the proofs of Proposition 2, and may also be of independent interest.
Proposition 3.
The following properties hold for the extended Bregman update mapping $\overline{T}_{\gamma}^{t}$:
  1. (i)
    (Well-posedness) The mapping $\overline{T}_{\gamma}^{t}(\bm{x})$ is well-defined and single-valued for all $\bm{x}\in{\cal X}$.
  2. (ii)
    (Boundedness) For any bounded sequence $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\cal X}$, the sequence $\{\overline{T}^{t}_{\gamma}(\bm{z}^{k})\}_{k\in{\mathbb{N}}}$ is also bounded.
  3. (iii)
    (Boundary coordinate consistency) Suppose that the sequence $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\cal X}$ converges to ${\overline{\bm{z}}}$ satisfying ${\cal I}(\bm{z}^{k})\equiv{\cal I}_{0}\subseteq[n]$, ${\cal B}(\bm{z}^{k})\equiv{\cal B}_{0}$, and $\overline{T}^{t}_{\gamma}(\bm{z}^{k})\to\overline{\bm{v}}\in{\cal X}$. Then, we have $\overline{\bm{v}}_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}$ and ${\cal B}(\overline{\bm{v}})={\cal B}(\overline{\bm{z}})$.
Proof.
(i) In view of Assumption 2 (iii), the function $G^{t}_{\gamma}(\ \cdot\ ;\bm{x})$ is strictly convex for any $\bm{x}\in{\cal X}$. To show the well-posedness of the extended Bregman update mapping, it suffices to verify that $G^{t}_{\gamma}(\ \cdot\ ;\bm{x})$ is level bounded. Lemma 3 establishes the coerciveness of $G^{t}_{\gamma}(\ \cdot\ ;\bm{x})$ for any fixed $\bm{x}\in{\cal X}$, and Assumption 2 (ii) ensures its lower semicontinuity. Hence, the level set $\{\bm{y}\in{\cal X}:G^{t}_{\gamma}(\bm{y};\bm{x})\leq c\}$ is compact for every $\bm{c}\in{\mathbb{R}}$, and the minimizer exists and is unique.
(ii) Let $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subset{\cal X}$ be a bounded sequence. WLOG, we may assume that $\bm{z}^{k}\to\overline{\bm{z}}$ by passing to a convergent subsequence if necessary. Since $G_{\gamma}^{t}(\bm{z}^{k};\bm{z}^{k})=\gamma(\bm{z}^{k};\bm{z}^{k})+g(\bm{z}^{k})$, and both $\gamma$ and $g$ are continuous by Assumption 2 (ii), we obtain:
$\lim_{k\to+\infty}G^{t}_{\gamma}(\bm{z}^{k};\bm{z}^{k})=\gamma(\overline{\bm{z}};\overline{\bm{z}})+g(\overline{\bm{z}})=f(\overline{\bm{z}})+g(\overline{\bm{z}})<+\infty.$
By the optimality of $\overline{T}^{t}_{\gamma}(\bm{z}^{k})$, we have $G^{t}_{\gamma}(\overline{T}^{t}_{\gamma}(\bm{z}^{k});\bm{z}^{k})\leq G^{t}_{\gamma}(\bm{z}^{k};\bm{z}^{k})$ and hence
$\limsup_{k\to+\infty}G^{t}_{\gamma}(\overline{T}^{t}_{\gamma}(\bm{z}^{k});\bm{z}^{k})\leq\lim_{k\to+\infty}G^{t}_{\gamma}(\bm{z}^{k};\bm{z}^{k})<+\infty.$
By Lemma 3, the sequence $\{\overline{T}^{t}_{\gamma}(\bm{z}^{k})\}_{k\in{\mathbb{N}}}$ is bounded.
(iii) Since ${\cal B}(\bm{z}^{k})\equiv{\cal B}_{0}$ and $\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{{\cal B}_{0}}=\bm{z}^{k}_{{\cal B}_{0}}$, taking limits gives $\overline{\bm{v}}_{{\cal B}_{0}}=\overline{\bm{z}}_{{\cal B}_{0}}$. Moreover, since $\bm{z}^{k}\to\overline{\bm{z}}$, we have ${\cal B}_{0}\subseteq{\cal B}(\overline{\bm{z}})$. It remains to verify coordinate equality on ${\cal B}(\overline{\bm{z}})\setminus{\cal B}_{0}\subseteq{\cal I}_{0}$.
We proceed by contradiction. Suppose that there exists $i_{0}\in{\cal I}_{0}\cap{\cal B}(\overline{\bm{z}})$ such that $\overline{v}_{i_{0}}\neq\overline{z}_{i_{0}}:=a$, and WLOG assume $a$ is the left endpoint of $\operatorname{cl}(\operatorname{dom}(\varphi))$. Since $z^{k}_{i_{0}}\to a$ and $\overline{T}_{\gamma}^{t}(z^{k})_{i_{0}}\to\overline{v}_{i_{0}}>a$, we have $z^{k}_{i_{0}}-\overline{T}_{\gamma}^{t}(\bm{z}^{k})_{i_{0}}\to a-\overline{v}_{i_{0}}<0$. Moreover, as $a$ is the left endpoint of $\operatorname{cl}(\operatorname{dom}(\varphi))$, it follows from Assumption 1 (iv) and Definition 1 (ii) that $\varphi^{\prime}(z^{k}_{i_{0}})\to\varphi^{\prime}(a)=-\infty$. Hence, we have
$\left(z^{k}_{i_{0}}-\overline{T}_{\gamma}^{t}(z^{k})_{i_{0}}\right)\left(\varphi^{\prime}(\overline{T}_{\gamma}^{t}(z^{k})_{i_{0}})-\varphi^{\prime}(z^{k}_{i_{0}})\right)\to-\infty.$
Subsequently, by the convexity of $\varphi$, we have $(z^{k}_{i}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i})(\varphi^{\prime}(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i})-\varphi^{\prime}(z^{k}_{i}))\leq 0$ for all $i\in[n]$, which leads to
$\sum_{i\in{\cal I}_{0}}\left(z^{k}_{i}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)\left(\varphi^{\prime}(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i})-\varphi^{\prime}(z^{k}_{i})\right)\rightarrow-\infty.$ (12)
On the other hand, we establish a finite lower bound for the same quantity in (12) by leveraging the optimality of $\overline{T}^{t}_{\gamma}(\bm{z}^{k})$, which will lead to a contradiction. To this end, we define a sequence of interpolation points
${\bm{z}}^{\theta,k}:=\theta\bm{z}^{k}+(1-\theta)\overline{T}^{t}_{\gamma}(\bm{z}^{k}),$ (13)
and a sequence of univariate functions as
${\phi}_{k}(\theta):=G^{t}_{\gamma}\left({\bm{z}}^{\theta,k};\bm{z}^{k}\right),\qquad\forall\ \theta\in[0,1].$
By the optimality of $\overline{T}_{\gamma}^{t}(\bm{z}^{k})$, we have $\phi_{k}(0)\leq\phi_{k}(\theta)$ for all $\theta\in[0,1]$, which implies $\frac{\phi_{k}(\theta)-\phi_{k}(0)}{\theta}\geq 0,$ for all $\theta\in(0,1]$. We next expand the difference quotient as follows:
$\frac{\phi_{k}(\theta)-\phi_{k}(0)}{\theta}$ $=\frac{\gamma\left(\bm{z}^{\theta,k};\bm{z}^{k}\right)-\gamma\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k});\bm{z}^{k}\right)}{\theta}+\frac{g\left(\bm{z}^{\theta,k}\right)-g\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})\right)}{\theta}$
$\quad+\frac{1}{t}\sum_{i\in{\cal I}_{0}}\frac{D_{\varphi}\left(z^{\theta,k}_{i},z^{k}_{i}\right)-D_{\varphi}\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z^{k}_{i}\right)}{\theta}.$
Letting $\theta\to 0_{+}$, and noting that $\bm{z}^{\theta,k}\to\overline{T}_{\gamma}^{t}(\bm{z}^{k})$, we obtain for all $i\in{\cal I}_{0}$, the following holds:
$\lim_{\theta\to 0_{+}}\frac{D_{\varphi}\left(z^{\theta,k}_{i},z^{k}_{i}\right)-D_{\varphi}\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z^{k}_{i}\right)}{\theta}$
$=$ $\lim_{\theta\to 0_{+}}\left\{\frac{\varphi\left(z^{\theta,k}_{i}\right)-\varphi\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)}{\theta}-\left\langle\varphi^{\prime}\left(z^{k}_{i}\right),\frac{z^{\theta,k}_{i}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}}{\theta}\right\rangle\right\}$
$=$ $\lim_{\theta\to 0_{+}}\left\{\frac{\varphi\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}+\theta\left(z_{i}^{k}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)\right)-\varphi\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)}{\theta}\right\}-\varphi^{\prime}(z^{k}_{i})\left(z^{k}_{i}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)$
$=$ $\left(z^{k}_{i}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)\left(\varphi^{\prime}\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)-\varphi^{\prime}(z^{k}_{i})\right),$
where the first equality follows from the definition of Bregman divergence, the second one follows from (13), and the last from the fact that $\varphi$ is continuous differentiable on ${\rm int}(\operatorname{dom}(\varphi))$. At boundary points, we use the extended directional derivative, i.e., $\varphi^{\prime}(a)=-\infty$ and $\varphi^{\prime}(c)=+\infty$.
We now substitute the above limit into the expansion of the difference quotient, which yields
$\limsup_{\theta\to 0_{+}}\frac{\phi_{k}(\theta)-\phi_{k}(0)}{\theta}$ $=\limsup_{\theta\to 0_{+}}\left[\frac{\gamma(\bm{z}^{\theta,k};\bm{z}^{k})-\gamma(\overline{T}_{\gamma}^{t}(\bm{z}^{k});\bm{z}^{k})}{\theta}+\frac{g(\bm{z}^{\theta,k})-g(\overline{T}_{\gamma}^{t}(\bm{z}^{k}))}{\theta}\right]$
$\quad+\frac{1}{t}\sum_{i\in\mathcal{I}_{0}}(z^{k}_{i}-\overline{T}_{\gamma}^{t}(\bm{z}^{k})_{i})\left(\varphi^{\prime}(\overline{T}_{\gamma}^{t}(\bm{z}^{k})_{i})-\varphi^{\prime}(z^{k}_{i})\right)\geq 0.$ (14)
Since the first two terms in the right-hand side of (5.1) are uniformly bounded, as ensured by the continuous differentiability of $\gamma(\cdot;\bm{z}^{k})$, the local Lipschitz continuity of $g$, and the boundedness of the sequences $\{\bm{z}^{k}\}_{k\in\mathbb{N}}$ and $\{\overline{T}^{t}_{\gamma}(\bm{z}^{k})\}_{k\in\mathbb{N}}$, we obtain
$\liminf_{k\to\infty}\sum_{i\in{\cal I}_{0}}\left(z^{k}_{i}-\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i}\right)\left(\varphi^{\prime}(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i})-\varphi^{\prime}(z^{k}_{i})\right)>-\infty,$
which contradicts (12). Therefore, we conclude that $\overline{\bm{v}}_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}$, which directly implies ${\cal B}(\overline{\bm{z}})\subseteq{\cal B}(\overline{\bm{v}})$.
By applying the same argument with the roles of $\overline{\bm{z}}$ and $\overline{\bm{v}}$ reversed, we similarly obtain $\overline{\bm{v}}_{{\cal B}(\overline{\bm{v}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{v}})}$, and hence ${\cal B}(\overline{\bm{v}})\subseteq{\cal B}(\overline{\bm{z}})$. It follows that ${\cal B}(\overline{\bm{v}})={\cal B}(\overline{\bm{z}})$, completing the proof.
Armed with the extended update mapping, we proceed to define the extended Bregman stationarity measure $\overline{R}^{t}_{\gamma}$ over the entire domain ${\cal X}$, whose well-definedness is ensured by that of $\overline{T}^{t}_{\gamma}$.
Definition 5 (Extended stationarity measure).
We define the extended Bregman stationarity measure $\overline{R}^{t}_{\gamma}(\bm{x}):{\cal X}\to{\mathbb{R}}$ as $\overline{R}^{t}_{\gamma}(\bm{x}):=\sum_{i\in{\cal I}(\bm{x})}D_{\varphi}(\overline{T}^{t}_{\gamma}(\bm{x})_{i},x_{i}).$
5.2 Proof of Proposition 1
Proof.
We first claim that $\overline{R}_{\gamma}^{t}(\overline{\bm{z}})=0$ if and only if $\overline{T}_{\gamma}^{t}(\overline{\bm{z}})=\overline{\bm{z}}$. By definition of $\overline{R}_{\gamma}^{t}$, we have:
$\overline{R}_{\gamma}^{t}(\overline{\bm{z}})=0\quad\Longleftrightarrow\quad\overline{T}_{\gamma}^{t}(\overline{\bm{z}})_{i}=\overline{z}_{i}\quad\text{for all }i\in{\cal I}(\overline{\bm{z}}).$
On the other hand, by construction of $\overline{T}_{\gamma}^{t}$, we always have $\overline{T}^{t}_{\gamma}(\overline{\bm{z}})_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}$ due to the hard constraint imposed on the boundary coordinates. Therefore, $\overline{T}_{\gamma}^{t}(\overline{\bm{z}})=\overline{\bm{z}}$ if and only if $\overline{R}_{\gamma}^{t}(\overline{\bm{z}})=0$.
It remains to show that $\overline{T}_{\gamma}^{t}(\overline{\bm{z}})=\overline{\bm{z}}$ if and only if there exists a subgradient $\bm{p}\in\partial F(\overline{\bm{z}})$ satisfying $\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}$. From the definition of the extended Bregman update mapping, we know that $\overline{T}_{\gamma}^{t}(\overline{\bm{z}})=\overline{\bm{z}}$ if and only if $\overline{\bm{z}}\in\operatorname*{argmin}_{\bm{y}\in{\cal X}}G_{\gamma}^{t}(\bm{y};\overline{\bm{z}})$. By the convexity of $G_{\gamma}^{t}(\cdot;\overline{\bm{z}})$, this is equivalent to the first-order optimality condition: ${\mathbf{0}}\in\partial G_{\gamma}^{t}(\overline{\bm{z}};\overline{\bm{z}}).$ According to Assumption 2 (ii) and (Rockafellar and Wets, 2009, Corollary 10.9), we have:
$\partial G^{t}_{\gamma}(\bm{y};\overline{\bm{z}})\mid_{\bm{y}=\overline{\bm{z}}}=\nabla f(\overline{\bm{z}})+\partial g(\overline{\bm{z}})+\partial\delta_{\bm{y}_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}}(\bm{y})\mid_{\bm{y}=\overline{\bm{z}}}.$ (15)
The last term corresponds to the subdifferential of the indicator function enforcing the constraint $\bm{y}_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}$, and satisfies: $\partial\delta_{\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}}(\overline{\bm{z}})=\operatorname{span}\{\bm{e}_{b}:b\in{\cal B}(\overline{\bm{z}})\}.$ Therefore, $\mathbf{0}\in\partial G_{\gamma}^{t}(\overline{\bm{z}};\overline{\bm{z}})$ if and only if there exists
$\bm{p}\in\nabla f(\overline{\bm{z}})+\partial g(\overline{\bm{z}})=\partial F(\overline{\bm{z}})$
such that $\bm{p}\in\operatorname{span}\{\bm{e}_{b}:b\in{\cal B}(\overline{\bm{z}})\}$, i.e., $\bm{p}_{{\cal I}(\overline{\bm{z}})}={\mathbf{0}}$. We complete the proof.
5.3 Proof of Proposition 2
Proof.
Based on the definition of $\overline{R}_{\gamma}^{t}$, our first goal is to establish the continuity of the extended Bregman update mapping. Specifically, we show that for any sequence $\{\bm{z}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\cal X}$ converging to $\overline{\bm{z}}\in{\cal X}$, it holds that
$\lim\limits_{k\to\infty}\overline{T}^{t}_{\gamma}(\bm{z}^{k})=\overline{T}^{t}_{\gamma}(\overline{\bm{z}}).$
By Proposition 3 (ii), the sequence $\{\overline{T}_{\gamma}^{t}(\bm{z}^{k})\}$ is bounded. We consider an arbitrary convergent subsequence and denote its limit by $\overline{\bm{v}}\in{\cal X}$. Moreover, since $\mathcal{B}(\bm{z}^{k})\subseteq[n]$ only takes values in a finite set, we may assume (by further subsequence selection if needed) that $\mathcal{B}(\bm{z}^{k})\equiv\mathcal{B}_{0}$ and $\mathcal{I}(\bm{z}^{k})\equiv\mathcal{I}_{0}:=[n]\setminus\mathcal{B}_{0}$. In the sequel, we will show that $\overline{\bm{v}}=\overline{T}_{\gamma}^{t}(\overline{\bm{z}})$. This implies that all convergent subsequences have the same limit, and hence the full sequence $\overline{T}_{\gamma}^{t}(\bm{z}^{k})$ converges to $\overline{T}_{\gamma}^{t}(\overline{\bm{z}})$.
Due to the strict convexity of $G^{t}_{\gamma}(\cdot;\overline{\bm{z}})$, as assumed in Assumption 2 (iii), and the optimality condition $\overline{T}^{t}_{\gamma}(\overline{\bm{z}})=\operatorname*{argmin}_{\bm{y}\in{\cal X}}\ G^{t}_{\gamma}(\bm{y};\overline{\bm{z}})$, it suffices to show
$G^{t}_{\gamma}(\overline{\bm{v}};\overline{\bm{z}})=G^{t}_{\gamma}\left(\overline{T}^{t}_{\gamma}(\overline{\bm{z}});\overline{\bm{z}}\right).$
Since the minimizer $\overline{T}^{t}_{\gamma}(\overline{\bm{z}})$ ensures $G^{t}_{\gamma}(\overline{\bm{v}};\overline{\bm{z}})\geq G^{t}_{\gamma}(\overline{T}^{t}_{\gamma}(\overline{\bm{z}});\overline{\bm{z}})$, it suffices to establish the reverse inequality:
$G^{t}_{\gamma}(\overline{\bm{v}};\overline{\bm{z}})\leq G^{t}_{\gamma}\left(\overline{T}^{t}_{\gamma}(\overline{\bm{z}});\overline{\bm{z}}\right).$ (16)
By Proposition 3 (iii), we have $\overline{\bm{v}}_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}$ and ${\cal B}(\overline{\bm{v}})={\cal B}(\overline{\bm{z}})$. Therefore, by the definition of $G^{t}_{\gamma}$ we can write
$G^{t}_{\gamma}(\overline{\bm{v}};\overline{\bm{z}})=\gamma(\overline{\bm{v}};\overline{\bm{z}})+g(\overline{\bm{v}})+\sum_{i\in{\cal I}(\overline{\bm{z}})}D_{\varphi}(\overline{v}_{i},\overline{z}_{i}).$
By the continuity of $\gamma$, $g$, and $D_{\varphi}$, together with the convergences $\bm{z}^{k}\to\overline{\bm{z}}$, $\overline{T}^{t}_{\gamma}(\bm{z}^{k})\to\overline{\bm{v}}$, we have
$G^{t}_{\gamma}(\overline{\bm{v}};\overline{\bm{z}})=\lim_{k\to\infty}\gamma\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k});\bm{z}^{k}\right)+g\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})\right)+\sum_{i\in{\cal I}(\overline{\bm{z}})}D_{\varphi}\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z^{k}_{i}\right).$
Since $\bm{z}^{k}\to\overline{\bm{z}}$, we have ${\cal I}(\bm{z}^{k})={\cal I}_{0}\supseteq{\cal I}(\overline{\bm{z}})$, which yields
$G^{t}_{\gamma}(\overline{\bm{v}};\overline{\bm{z}})$ $\leq\lim_{k\to\infty}\gamma\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k});\bm{z}^{k}\right)+g\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})\right)+\sum_{i\in{\cal I}_{0}}D_{\varphi}\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z^{k}_{i}\right)$ (17)
$=\lim_{k\to\infty}\min_{\bm{y}\in{\cal X}}\ G^{t}_{\gamma}(\bm{y};\bm{z}^{k}),$
where the equality follows from the definitions of $G^{t}_{\gamma}$ and $\overline{T}^{t}_{\gamma}$. To complete the proof of the reverse inequality (16), it remains to show that
$\lim_{k\to\infty}\min_{\bm{y}\in{\cal X}}G^{t}_{\gamma}(\bm{y};\bm{z}^{k})\leq G^{t}_{\gamma}(\overline{T}^{t}_{\gamma}(\overline{\bm{z}});\overline{\bm{z}}).$ (18)
To this end, we define constraint sets ${\cal C}^{k}$ by
${\cal C}^{k}={\cal X}\cap\left\{\bm{y}:\bm{y}_{{\cal B}(\overline{\bm{z}})}=\bm{z}^{k}_{{\cal B}(\overline{\bm{z}})}\right\},\qquad\forall~{}k\in{\mathbb{N}}.$
As $\bm{z}^{k}\in{\cal C}^{k}$, it follows that $\mathcal{C}^{k}$ is nonempty. Thanks to Rockafellar and Wets (2009, Exercise 4.33) and Rockafellar and Wets (2009, Theorem 4.32 (b)), the following set limit holds in the sense of Painlevé–Kuratowski convergence:
$\lim_{k\to\infty}{\cal C}^{k}=\overline{{\cal C}}:={\cal X}\cap\lim_{k\to\infty}\left\{\bm{y}:\bm{y}_{{\cal B}(\overline{\bm{z}})}={\bm{z}}^{k}_{{\cal B}(\overline{\bm{z}})}\right\}={\cal X}\cap\left\{\bm{y}:\bm{y}_{{\cal B}(\overline{\bm{z}})}=\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}\right\}.$ (19)
Moreover, due to Proposition 3 (iii), we know that $\overline{\bm{z}}_{{\cal B}(\overline{\bm{z}})}=\overline{T}^{t}_{\gamma}(\overline{\bm{z}})_{{\cal B}(\overline{\bm{z}})}$ and then $\overline{T}^{t}_{\gamma}(\overline{\bm{z}})\in\overline{{\cal C}}$. The set convergence (19) yields that there exists a sequence $\{\bm{y}^{k}\}_{k\in{\mathbb{N}}}$ with $\bm{y}^{k}\in{\cal C}^{k}\subseteq{\cal X}$ converging to $\overline{T}^{t}_{\gamma}(\overline{\bm{z}})$. Then, we have
$\lim_{k\to\infty}\min_{\bm{y}\in{\cal X}}G^{t}_{\gamma}(\bm{y};\bm{z}^{k})$ $\leq\lim_{k\to\infty}\ G^{t}_{\gamma}(\bm{y}^{k};\bm{z}^{k})=\lim_{k\to\infty}\gamma\left(\bm{y}^{k};\bm{z}^{k}\right)+g\left(\bm{y}^{k}\right)+\sum_{i\in{\cal I}_{0}}D_{\varphi}\left(y^{k}_{i},z^{k}_{i}\right)$
$=\lim_{k\to\infty}\gamma\left(\bm{y}^{k};\bm{z}^{k}\right)+g\left(\bm{y}^{k}\right)+\sum_{i\in{\cal I}(\overline{\bm{z}})}D_{\varphi}\left(y^{k}_{i},z^{k}_{i}\right)=G^{t}_{\gamma}\left(\overline{T}^{t}_{\gamma}(\overline{\bm{z}});\overline{\bm{z}}\right),$
where the second equality follows from the facts that ${\cal I}_{0}\supseteq{\cal I}(\overline{\bm{z}})$ and $\bm{y}^{k}_{{\cal B}(\overline{\bm{z}})}=\bm{z}^{k}_{{\cal B}(\overline{\bm{z}})}$, and the final equality follows from the continuity of $\gamma$, $g$, and $D_{\varphi}$. This completes the proof of (18), and we thus conclude that the mapping $\overline{T}^{t}_{\gamma}$ is continuous.
We are now ready to prove the continuity of $\overline{R}^{t}_{\gamma}$. WLOG, we may assume that $\mathcal{I}(\bm{z}^{k})\equiv\mathcal{I}_{0}\subseteq[n]$ for all $k\in{\mathbb{N}}$. Since $\bm{z}^{k}\to\overline{\bm{z}}$, it follows that $\mathcal{I}(\overline{\bm{z}})\subseteq\mathcal{I}_{0}$. We proceed by analyzing the two index sets $\mathcal{I}(\overline{\bm{z}})$ and $\mathcal{I}_{0}\setminus\mathcal{I}(\overline{\bm{z}})$ separately.
  1. (i)
    For $i\in{\cal I}(\overline{\bm{z}})$, the continuity of $\overline{T}^{t}_{\gamma}$ and the convergence $z^{k}_{i}\to\overline{z}_{i}\in{\rm{int}}(\operatorname{dom}(\varphi))$ directly imply
    $D_{\varphi}\left(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z_{i}^{k}\right)\to D_{\varphi}\left(\overline{T}^{t}_{\gamma}(\overline{\bm{z}})_{i},\overline{z}_{i}\right),\qquad\forall\ i\in{\cal I}(\overline{\bm{z}}).$
  2. (ii)
    For $i\in\mathcal{I}_{0}\setminus\mathcal{I}(\overline{\bm{z}})$, from the proof establishing the continuity of $\overline{T}^{t}_{\gamma}$, we know that the inequalities in (17) and (18) in fact hold with equality. Then, we have
    $G^{t}_{\gamma}(\overline{T}^{t}_{\gamma}(\overline{\bm{z}});\overline{\bm{z}})=\lim_{k\to\infty}\gamma(\overline{T}^{t}_{\gamma}(\bm{z}^{k});\bm{z}^{k})+g(\overline{T}^{t}_{\gamma}(\bm{z}^{k}))+\sum_{i\in\mathcal{I}_{0}}D_{\varphi}(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z^{k}_{i}).$
    Combining this together with the definition of $G^{t}_{\gamma}$, we have
    $D_{\varphi}(\overline{T}^{t}_{\gamma}(\bm{z}^{k})_{i},z^{k}_{i})\to 0,\qquad\forall\,i\in\mathcal{I}_{0}\setminus\mathcal{I}(\overline{\bm{z}}).$
Combining the two cases, we conclude that $\overline{R}^{t}_{\gamma}(\bm{z}^{k})\to\overline{R}^{t}_{\gamma}(\overline{\bm{z}})$, thereby establishing the continuity of $\overline{R}^{t}_{\gamma}$. ∎
6 Closing Remarks
This paper uncovers a fundamental limitation of BPs: BPs can stall near spurious stationary points due to degeneracies in Bregman geometry. We show that such points arise generically, mislead standard residual-based stationarity measure, and cause finite-time stagnation without reliable convergence guarantees. These findings challenge core assumptions of existing Bregman-based methods and underscore the need for new algorithmic designs and theoretical frameworks. An important direction for future research is to develop heuristic safeguards or algorithmic modifications that can prevent finite-time trapping near spurious stationary points. Designing such strategies remains an open and practically relevant question.
Acknowledgment
Jiajin Li thanks Prof. Heinz H. Bauschke, Prof. Joseph Paat and Prof. Yinyu Ye for their helpful discussions. Jiajin Li was supported by a Natural Sciences and Engineering Research Council of Canada Discovery Grant GR034865.
Appendix A Verification of Assumption 2 (iv)
When $\operatorname{dom}(\varphi)$ is open—for example, when $\varphi(x)=\tfrac{1}{x}$—the condition (3) typically fails. To ensure the well-posedness of BPs in such cases, it is necessary to invoke the compactness of ${\cal X}$, as required in Assumption 2 (iv). Such a supplementary condition is also standard in the classical Bregman literature; see, for instance, condition (i) in Bauschke et al. [2017, Lemma 2]. In the following lemma, we focus on the case where $\operatorname{dom}(\varphi)$ is closed and examine how Assumption 2 (iv) relates to existing conditions in the literature.
Lemma 4.
Suppose that Assumption 1 holds and $\text{dom}(\varphi)$ is closed. The following statements hold:
  1. (i)
    If the surrogate model takes the form $\gamma(\bm{y};\bm{x})=f(\bm{x})+\nabla f(\bm{x})^{\top}(\bm{y}-\bm{x})$ and $h+tg$ is supercoercive for all $t\in(0,\overline{t}]$ for some $\overline{t}>0$, then Assumption 2 (iv) is satisfied.
  2. (ii)
    If the surrogate model takes the form $\gamma(\bm{y};\bm{x})=f(\bm{y})$ and $h+tF$ is convex for all $t\in(0,\overline{t}]$ for some $\overline{t}>0$, then Assumption 2 (iv) is satisfied.
  3. (iii)
    If the surrogate model takes the form $\gamma(\bm{y};\bm{x})=f(\bm{x})+\nabla f(\bm{x})^{\top}(\bm{y}-\bm{x})+\frac{1}{2}(\bm{y}-\bm{x})^{\top}\nabla^{2}f(\bm{x})(\bm{y}-\bm{x})$, $h+tg$ is supercoercive for all for all $t\in(0,\overline{t}]$ for some $\overline{t}>0$, and $f$ is a convex function, then Assumption 2 (iv) is satisfied.
Remark 3.
As shown in our proof of Lemma 4 (ii), the condition that $h+tg$ is supercoercive for all $t\in(0,\overline{t}]$ can be removed as we know $h+tg$ is convex from Assumption 1.
Remark 4.
The assumptions made in Lemma 4 are consistent with standard practices in the literature: (i) For the Bregman proximal gradient method with surrogate model $\gamma(\bm{y};\bm{x})=f(\bm{x})+\nabla f(\bm{x})^{\top}(\bm{y}-\bm{x})$, the supercoercivity of $h+tg$ is a common assumption; see, e.g., [Bauschke et al., 2017, Lemma 2] and [Bolte et al., 2018, Assumption B]. (ii) The convexity of $h+tF$ is a standard assumption in the analysis of Bregman proximal point methods; see, e.g., [Zhang and He, 2018, Assumption 3.1 (ii)] and [Chen and Teboulle, 1993]. (iii) The convexity condition on $f$ in Lemma 4 (iii) is also assumed in Doikov and Nesterov [2023].
Proof.
(i): To verify Assumption 2 (iv), we prove the stronger statement:
$\lim\limits_{k\to\infty}\frac{\gamma(\bm{y}^{k};\bm{z}^{k})+g(\bm{y}^{k})+\frac{1}{t}D_{h}(\bm{y}^{k},\bm{z}^{k})}{\|\bm{y}_{k}\|}=+\infty.$
Since $\|\bm{y}^{k}\|\to\infty$, $\bm{z}^{k}\to\overline{\bm{z}}$, and $\nabla f$ is continuous, we have
$\lim\limits_{k\to\infty}\frac{\gamma(\bm{y}^{k};\bm{z}^{k})}{\|\bm{y}^{k}\|}=\nabla f(\overline{\bm{z}})^{\top}\lim\limits_{k\to\infty}\frac{\bm{y}^{k}}{\|\bm{y}^{k}\|}<+\infty.$
Therefore, it suffices to show that
$\lim\limits_{k\to\infty}\frac{g(\bm{y}^{k})+\frac{1}{t}D_{h}(\bm{y}^{k},\bm{z}^{k})}{\|\bm{y}^{k}\|}=+\infty.$
We now estimate the Bregman distance term $D_{h}(\bm{y}^{k},\bm{z}^{k})$, which decomposes coordinate-wise as
$\sum_{i\in{\cal I}(\overline{\bm{z}})}D_{\varphi}(y^{k}_{i},z^{k}_{i})+\sum_{b\in{\cal B}(\overline{\bm{z}})}D_{\varphi}(y^{k}_{b},z^{k}_{b}),$
where
${\cal B}(\overline{\bm{z}})=\{b\in[n]:\overline{z}_{b}\in\operatorname{bd}(\operatorname{dom}(\varphi))=\{a,c\}\}\text{ and }{\cal I}(\overline{\bm{z}})=\{i\in[n]:\overline{z}_{i}\in{\rm int}(\operatorname{dom}(\varphi))=(a,c)\}.$
We analyze the interior and boundary coordinates separately.
  1. (a)
    For $i\in{\cal I}(\overline{\bm{z}})$, we have $\lim\limits_{k\to\infty}|\varphi^{\prime}(z_{i}^{k})|=|\varphi^{\prime}(\overline{z}_{i})|<+\infty$ due to the continuous differentiability of $\varphi$ on ${\rm int}(\operatorname{dom}(\varphi))$. It follows that
    $\lim\limits_{k\to\infty}\frac{D_{\varphi}(y^{k}_{i},z^{k}_{i})}{\|\bm{y}^{k}\|}$ $\,=\lim\limits_{k\to\infty}\frac{\varphi(y_{i}^{k})-\varphi(z_{i}^{k})-\varphi^{\prime}(z_{i}^{k})(y_{i}^{k}-z_{i}^{k})}{\|\bm{y}^{k}\|},$
    $\,=\lim\limits_{k\to\infty}\frac{\varphi(y_{i}^{k})}{\|\bm{y}^{k}\|}-\varphi^{\prime}(\overline{z}_{i})\cdot\lim\limits_{k\to\infty}\frac{y_{i}^{k}}{\|\bm{y}^{k}\|},\quad\forall\ i\in{\cal I}(\overline{\bm{z}}).$ (20)
    The first equality follows from the definition of $D_{\varphi}$ and the second uses the convergence of $\bm{z}^{k}\to\overline{\bm{z}}$ and the finiteness of $\varphi(\overline{z}_{i})$ and $\varphi^{\prime}(\overline{z}_{i})$ for $i\in{\cal I}(\overline{\bm{z}})$.
  2. (b)
    For $b\in{\cal B}(\overline{\bm{z}})$, WLOG, we may assume that $\overline{z}_{b}=a$, which is finite since $\overline{\bm{z}}\in{\cal X}\subseteq{\mathbb{R}}^{n}$.
    (1) When $y_{b}^{k}\leq z_{b}^{0}$, we have $a\leq y_{b}^{k}\leq z_{b}^{0}$. By (2), we have $D_{\varphi}(y^{k}_{b},z^{0}_{b})\leq D_{\varphi}(a,z^{0}_{b})$, and thus
    $D_{\varphi}(y^{k}_{b},z^{k}_{b})\geq 0\geq D_{\varphi}(y^{k}_{b},z^{0}_{b})-D_{\varphi}(a,z^{0}_{b}).$
    (2) When $y_{b}^{k}>z_{b}^{0}$, we have $z_{b}^{k}\leq z_{b}^{0}<y_{b}^{k}$ for sufficiently large $k$ due to the convergence $\lim_{k\to\infty}z_{b}^{k}=\overline{z}_{b}=a$. Again using the same inequality (2), we get
    $D_{\varphi}(y^{k}_{b},z^{k}_{b})\geq D_{\varphi}(y^{k}_{b},z^{0}_{b}).$
    Note that $D_{\varphi}(a,z^{0}_{b})$ is finite, since $\overline{z}_{b}=a$ is finite and $\varphi$ is a closed function. By considering both cases, we obtain
    $\lim\limits_{k\to\infty}\frac{D_{\varphi}(y^{k}_{b},z^{k}_{b})}{\|\bm{y}^{k}\|}\geq\lim\limits_{k\to\infty}\frac{D_{\varphi}(y^{k}_{b},z^{0}_{b})}{\|\bm{y}^{k}\|}=\lim\limits_{k\to\infty}\frac{\varphi(y^{k}_{b})}{\|\bm{y}^{k}\|}-\varphi^{\prime}(z^{0}_{b})\lim\limits_{k\to\infty}\frac{y^{k}_{b}}{\|\bm{y}^{k}\|},\quad\forall~{}b\ \in{\cal B}(\overline{\bm{z}}).$ (21)
Combining ((a)) and (21), we obtain
$\lim\limits_{k\to\infty}\frac{\frac{1}{t}D_{h}(\bm{y}^{k},\bm{z}^{k})+g(\bm{y}^{k})}{\|\bm{y}^{k}\|}$
$\geq$ $\lim\limits_{k\to\infty}\frac{\frac{1}{t}h(\bm{y}^{k})+g(\bm{y}^{k})}{\|\bm{y}^{k}\|}-\frac{1}{t}\sum_{i\in{\cal I}(\overline{\bm{z}})}\varphi^{\prime}(\overline{z}_{i})\cdot\lim\limits_{k\to\infty}\frac{y_{i}^{k}}{\|\bm{y}^{k}\|}-\frac{1}{t}\sum_{b\in{\cal B}(\overline{\bm{z}})}\varphi^{\prime}(z_{b}^{0})\cdot\lim\limits_{k\to\infty}\frac{y_{b}^{k}}{\|\bm{y}^{k}\|}=+\infty,$
where the equality holds since $h+tg$ is supercoercive for all $t>0$, and the finiteness of $\varphi^{\prime}(z_{b}^{0})$ and $\varphi^{\prime}(\overline{z}_{i})$ for all $b\in{\cal B}(\overline{\bm{z}})$ and $i\in{\cal I}(\overline{\bm{z}})$. This completes the proof of (i).
(ii) We first show that $h+tF$ is supercoercive for all $t\in(0,\tfrac{\overline{t}}{2}]$, which follows from the convexity of $h+tF$ for $t\in(0,\overline{t}]$. Once this is established, the verification of Assumption 2 (iv) proceeds analogously to the proof of (i), with $g$ replaced by $F$.
By the convexity of $h+tF$ for $t\in(0,\overline{t}]$, we have for any $\bm{y}^{0}\in{\rm int}(\operatorname{dom}(h))\cap{\cal X}$ and any $t\in(0,\overline{t}]$,
$h(\bm{y})+tF(\bm{y})\geq h(\bm{y}^{0})+tF(\bm{y}^{0})+\left(\nabla h(\bm{y}^{0})+t\bm{v}^{0}\right)^{\top}(\bm{y}-\bm{y}^{0}),$
where $\bm{v}^{0}\in\partial F(\bm{y}^{0})$. It then follows that for any sequence $\{\bm{y}^{k}\}_{k\in{\mathbb{N}}}\subseteq{\rm int}(\operatorname{dom}(h))\cap{\cal X}$ with $\|\bm{y}^{k}\|\to\infty$ and any $t\in(0,\overline{t}]$, we have
$h(\bm{y}^{k})+\frac{t}{2}F(\bm{y}^{k})$ $=\frac{1}{2}h(\bm{y}^{k})+\frac{1}{2}\left(h(\bm{y}^{k})+tF(\bm{y}^{k})\right)$
$\geq\frac{1}{2}h(\bm{y}^{k})+\frac{1}{2}\left(h(\bm{y}^{0})+tF(\bm{y}^{0})+\left(\nabla h(\bm{y}^{0})+t\bm{v}^{0}\right)^{\top}(\bm{y}^{k}-\bm{y}^{0})\right).$
Using the lower bound above, we obtain
$\lim_{k\to\infty}\frac{h(\bm{y}^{k})+\frac{t}{2}F(\bm{y}^{k})}{\|\bm{y}^{k}\|}\geq\lim_{k\to\infty}\frac{1}{2}\frac{h(\bm{y}^{k})}{\|\bm{y}^{k}\|}+\frac{1}{2}\left(\nabla h(\bm{y}^{0})+t\bm{v}^{0}\right)^{\top}\lim_{k\to\infty}\frac{\bm{y}^{k}}{\|\bm{y}^{k}\|}.$
Hence, to establish supercoercivity, it suffices to show $\lim_{k\to\infty}\frac{h(\bm{y}^{k})}{\|\bm{y}^{k}\|}=+\infty$.
WLOG, we may assume the coordinate with the largest magnitude is index $i_{0}$, i.e., $i_{0}\equiv\operatorname*{argmax}_{i\in[n]}\{|y^{k}_{i}|:i\in[n]\}$ for all $k\in{\mathbb{N}}$. Then, $|y^{k}_{i_{0}}|\to+\infty$ and $|y_{i_{0}}^{k}|\geq\frac{\|\bm{y}^{k}\|}{n}$.
By the convexity of $\varphi$, for any $i\neq i_{0}$, we have
$\lim_{k\to\infty}\frac{\varphi(y^{k}_{i})}{\|\bm{y}^{k}\|}\geq\lim_{k\to\infty}\frac{\varphi(y^{0}_{i})+\varphi^{\prime}(y^{0}_{i})(y^{k}_{i}-y^{0}_{i})}{\|\bm{y}^{k}\|}=\lim_{k\to\infty}\varphi^{\prime}(y^{0}_{i})\frac{y^{k}_{i}}{\|\bm{y}^{k}\|}>-\infty,$
where the equality follows from $\|\bm{y}^{k}\|\to+\infty$. For $i=i_{0}$, we have
$\lim_{k\to\infty}\frac{\varphi(y^{k}_{i_{0}})}{\|\bm{y}^{k}\|}$ $\geq\lim_{k\to\infty}\frac{\varphi\left(\frac{y^{k}_{i_{0}}+y^{0}_{i_{0}}}{2}\right)+\frac{1}{2}\varphi^{\prime}\left(\frac{y^{k}_{i_{0}}+y^{0}_{i_{0}}}{2}\right)(y^{k}_{i_{0}}-y^{0}_{i_{0}})}{\|\bm{y}^{k}\|}$
$\geq\lim_{k\to\infty}\frac{\varphi(y^{0}_{i_{0}})+\frac{1}{2}\varphi^{\prime}(y^{0}_{i_{0}})(y^{k}_{i_{0}}-y^{0}_{i_{0}})+\frac{1}{2}\varphi^{\prime}\left(\frac{y^{k}_{i_{0}}+y^{0}_{i_{0}}}{2}\right)(y^{k}_{i_{0}}-y^{0}_{i_{0}})}{\|\bm{y}^{k}\|}$
$=\frac{1}{2}\lim_{k\to\infty}\frac{\varphi^{\prime}\left(\frac{y^{k}_{i_{0}}+y^{0}_{i_{0}}}{2}\right)y^{k}_{i_{0}}}{\|\bm{y}^{k}\|}+\frac{1}{2}\lim_{k\to\infty}\frac{\varphi^{\prime}(y^{0}_{i_{0}})y^{k}_{i_{0}}}{\|\bm{y}^{k}\|}$
$\geq\frac{1}{2n}\lim_{k\to\infty}\left|\varphi^{\prime}\left(\frac{y^{k}_{i_{0}}+y^{0}_{i_{0}}}{2}\right)\right|-\frac{1}{2}\left|\varphi^{\prime}(y^{0}_{i_{0}})\right|=+\infty,$
where the first and second inequalities follow from the convexity of $\varphi$; the third one uses the bound $\|\bm{y}^{k}\|\geq|y_{i_{0}}^{k}|\geq\frac{\|\bm{y}^{k}\|}{n}$; and the final equality follows from Assumption 1 (iv) and Definition 1 (ii).
(iii): Given the convexity of $f$, we have $(\bm{y}-\bm{x})^{\top}\nabla^{2}f(\bm{x})(\bm{y}-\bm{x})\geq 0$ for all $\bm{y}\in{\mathbb{R}}^{n}$. Hence, it suffices to show that
$\lim\limits_{k\to\infty}{f(\bm{z}^{k})+\nabla f(\bm{z}^{k})^{T}(\bm{y}^{k}-\bm{z}^{k})+g(\bm{y}^{k})+\frac{1}{t}D_{h}(\bm{y}^{k},\bm{z}^{k})}=+\infty,$
which is exactly the setting in (i). We finish the proof. ∎
Appendix B Missing Proofs for Example 3
The omitted technical details for Example 3 are provided here. We first verify that the function defined in (11) satisfies the assumptions stated in Example 3.
Fact 1.
Suppose that $\alpha\in(0,0.1]$ and consider the Burg entropy kernel $\varphi(x)=-\log(x)$. The function $f_{\alpha}:{\mathbb{R}}^{2}\rightarrow{\mathbb{R}}$ defined in Example 3, with $\phi_{\alpha}$ given by (11), satisfies the following properties:
  1. (i)
    $\phi_{\alpha}$ is continuously differentiable on the open domain $(0,1)\times(0,1)$;
  2. (ii)
    $\phi_{\alpha}(x)=2(x-\alpha)$ for $x\geq a$, $\phi_{\alpha}(x)\leq 0$ for $x\leq\alpha$, and $\phi_{\alpha}(0)\leq-1$;
  3. (iii)
    $\phi^{\prime}_{\alpha}(x)\geq 1$ for $x\in[0,1]$;
  4. (iv)
    $f_{\alpha}+h$ and $-f_{\alpha}+h$ are convex on $(0,1]\times(0,1]$.
Proof.
(i): Establishing differentiability over the entire open interval reduces to verifying both continuity and differentiability of $\phi_{\alpha}$ at the junction points where the functional form changes. These points are:
$x=\alpha,\quad x=\alpha\exp\left(-\frac{1}{\alpha}\right),\quad\text{and}\quad x=\frac{1}{2}\alpha\exp\left(-\frac{1}{\alpha}\right).$
As a representative case, we verify continuity and differentiability at $x=\alpha$:
  • For $x\geq\alpha$, we have $\phi_{\alpha}(x)=2x-2\alpha$, which gives $\phi_{\alpha}(\alpha^{+})=0$ and $\phi_{\alpha}^{\prime}(\alpha^{+})=2$.
  • For $x\in\left[\alpha\exp\left(-\tfrac{1}{\alpha}\right),\alpha\right]$, we have $\phi_{\alpha}(x)=x+\alpha\log\left(\tfrac{x}{\alpha}\right)-\alpha$. Hence, we have
    $\phi_{\alpha}(\alpha^{-})=\alpha+\alpha\log(1)-\alpha=0,\quad\text{and }\quad\phi_{\alpha}^{\prime}(\alpha^{-})=1+\frac{\alpha}{x}\Big{|}_{x=\alpha}=2.$
Thus, $\phi_{\alpha}$ is continuously differentiable at $x=\alpha$. Similar computations at the other two junction points confirm that $\phi_{\alpha}$ and its derivative are continuous at those points as well.
(ii): This follows directly from the definition of $\phi_{\alpha}(x)$ in (11):
  • For $x\geq\alpha$, $\phi_{\alpha}(x)=2(x-\alpha)$ by construction.
  • For $x\leq\alpha$, every piece of $\phi_{\alpha}(x)$ is nonpositive.
  • In particular, $\phi_{\alpha}(0)=\alpha(\log 2-2)-1\leq-1$ for all $\alpha\in(0,0.1]$.
(iii): We verify that $\phi_{\alpha}^{\prime}(x)\geq 1$ for all $x\in[0,1]$ by examining the derivative on each interval specified in (11).
  • For $x\in[\alpha,1]$, we have $\phi_{\alpha}^{\prime}(x)=2$.
  • For $x\in\left[\alpha\exp\left(-\tfrac{1}{\alpha}\right),\alpha\right]$, we have $\phi_{\alpha}^{\prime}(x)=1+\frac{\alpha}{x}\geq 2$.
  • For $x\in\left[\tfrac{1}{2}\alpha\exp\left(-\tfrac{1}{\alpha}\right),\alpha\exp\left(-\tfrac{1}{\alpha}\right)\right]$, we have $\phi_{\alpha}^{\prime}(x)=1-\frac{\alpha}{x}+2\exp\left(\tfrac{1}{\alpha}\right)\geq 1$.
  • For $x\in\left[0,\tfrac{1}{2}\alpha\exp\left(-\tfrac{1}{\alpha}\right)\right]$, we have $\phi_{\alpha}^{\prime}(x)=1$.
Therefore, $\phi_{\alpha}^{\prime}(x)\geq 1$ for all $x\in[0,1]$.
(iv): We begin by verifying that, on each piece of $f_{\alpha}$, the Hessians of $h+f_{\alpha}$ and $h-f_{\alpha}$ are positive semidefinite. To this end, we examine each case of $\phi_{\alpha}(x_{1})$ piece by piece:
  • For $x_{1}\in[\alpha,1]$, we have
    $\frac{1}{x_{1}^{2}}-|\nabla^{2}_{11}f_{\alpha}(\bm{x})|=\frac{1}{x_{1}^{2}}\geq|\nabla_{12}f_{\alpha}(\bm{x})|=1,\text{ and }\frac{1}{x_{2}^{2}}-|\nabla^{2}_{22}f_{\alpha}(\bm{x})|=\frac{1}{x_{2}^{2}}\geq|\nabla_{21}f_{\alpha}(\bm{x})|=1.$
    This implies that $\nabla^{2}h+\nabla^{2}f_{\alpha}$ and $\nabla^{2}h-\nabla^{2}f_{\alpha}$ are diagonally dominant with strictly positive diagonal entries, and are therefore positive semidefinite.
  • For $x_{1}\in\left[\alpha\exp\left(-\tfrac{1}{\alpha}\right),\alpha\right]$, we have
    $\frac{1}{x_{1}^{2}}-|\nabla_{11}^{2}f_{\alpha}(\bm{x})|=\frac{1}{x_{1}^{2}}-\frac{\alpha}{2}\cdot\frac{1}{x_{1}^{2}}(x_{2}+0.05)\geq\frac{0.9}{x_{1}^{2}}>|\nabla_{12}f_{\alpha}(\bm{x})|=\frac{1}{2}\left(1+\frac{\alpha}{x_{1}}\right),$
    where the last inequality follows from: (i) the function $x\mapsto\tfrac{0.9}{x^{2}}-\frac{1}{2}(1+\tfrac{\alpha}{x})$ monotonically decreases on $[0,\alpha]$; and (ii) its lower bound satisfies $\tfrac{0.9}{\alpha^{2}}-\frac{1}{2}(1+\tfrac{\alpha}{\alpha})\geq 0$ by $\alpha\leq 0.1$.
    Similarly, using $\frac{1}{x_{1}^{2}}-|\nabla_{11}^{2}f_{\alpha}(\bm{x})|\geq\frac{0.9}{x_{1}^{2}}$ and $\frac{1}{x_{2}^{2}}-|\nabla_{22}^{2}f_{\alpha}(\bm{x})|=\frac{1}{x_{2}^{2}}\geq 1$, we have
    $\left(\frac{1}{x_{1}^{2}}-|\nabla_{11}^{2}f_{\alpha}(\bm{x})|\right)\left(\frac{1}{x_{2}^{2}}-|\nabla_{22}^{2}f_{\alpha}(\bm{x})|\right)\geq\frac{0.9}{x_{1}^{2}}\geq|\nabla_{12}f_{\alpha}(\bm{x})|\cdot|\nabla_{21}f_{\alpha}(\bm{x})|=\frac{1}{4}\left(1+\frac{\alpha}{x_{1}}\right)^{2}.$
    Hence, all leading principal minors of $\nabla^{2}h+\nabla^{2}f_{\alpha}$ and $\nabla^{2}h-\nabla^{2}f_{\alpha}$ are positive, which implies that both matrices are positive definite.
  • For $x_{1}\in\left[\tfrac{1}{2}\alpha\exp\left(-\tfrac{1}{\alpha}\right),\alpha\exp\left(-\tfrac{1}{\alpha}\right)\right]$, we have
    $\frac{1}{x_{1}^{2}}-|\nabla^{2}_{11}f_{\alpha}(\bm{x})|=\frac{1}{x_{1}^{2}}-\frac{\alpha}{2}\cdot\frac{1}{x^{2}_{1}}(x_{2}+0.05)\geq\frac{0.9}{x_{1}^{2}}\geq|\nabla_{12}f_{\alpha}(\bm{x})|=\frac{1}{2}\left(1-\frac{\alpha}{x_{1}}+2\exp\left(\frac{1}{\alpha}\right)\right),$
    where the last inequality follows from: (i) the function $x\mapsto\frac{0.9}{x^{2}}+\frac{\alpha}{2x}-\frac{1}{2}-\exp(\tfrac{1}{\alpha})$ monotonically decreasing on $(0,1]$; (ii) its lower bound satisfies $\frac{0.9}{\alpha^{2}}\exp\left(\tfrac{2}{\alpha}\right)-\frac{1}{2}-\frac{1}{2}\exp(\frac{1}{\alpha})\geq 0$ by $\alpha\leq 0.1$.
    Moreover, using the estimates
    $\frac{1}{x_{1}^{2}}-|\nabla_{11}^{2}f_{\alpha}(\bm{x})|\geq\frac{0.9}{x_{1}^{2}};\qquad\frac{1}{x_{2}^{2}}-|\nabla_{22}^{2}f_{\alpha}(\bm{x})|=\frac{1}{x_{2}^{2}}\geq 1;$
    $|\nabla_{12}f_{\alpha}(\bm{x})|=|\nabla_{21}f_{\alpha}(\bm{x})|=\frac{1}{2}\left(1-\frac{\alpha}{x_{1}}+2\exp\left(\frac{1}{\alpha}\right)\right)\leq\frac{1}{2}+\exp\left(\frac{1}{\alpha}\right)\leq 2\exp\left(\frac{1}{\alpha}\right),$
    we have
    $\left(\frac{1}{x_{1}^{2}}-|\nabla_{11}^{2}f_{\alpha}(\bm{x})|\right)\left(\frac{1}{x_{2}^{2}}-|\nabla_{22}^{2}f_{\alpha}(\bm{x})|\right)\geq\frac{0.9}{x_{1}^{2}}>\left(2\exp\left(\frac{1}{\alpha}\right)\right)^{2}\geq|\nabla_{12}f_{\alpha}(\bm{x})|\cdot|\nabla_{21}f_{\alpha}(\bm{x})|.$
    Here, the second inequality follows from $x_{1}\in\left[\tfrac{1}{2}\alpha\exp\left(-\tfrac{1}{\alpha}\right),\alpha\exp\left(-\tfrac{1}{\alpha}\right)\right]$ and $\alpha\leq 0.1$, which together imply that
    $\frac{0.9}{x_{1}^{2}}\geq\frac{0.9\exp\left(\frac{2}{\alpha}\right)}{\alpha^{2}}\geq 90\exp\left(\frac{2}{\alpha}\right)>\left(2\exp\left(\frac{1}{\alpha}\right)\right)^{2}.$
    This confirms that both $\nabla^{2}h+\nabla^{2}f_{\alpha}$ and $\nabla^{2}h-\nabla^{2}f_{\alpha}$ have positive leading principal minors, and thus are positive definite.
  • For $x_{1}\in\left[0,\tfrac{1}{2}\alpha\exp\left(-\tfrac{1}{\alpha}\right)\right]$, the same reasoning as in the first case applies.
Consequently, both $h+f_{\alpha}$ and $h-f_{\alpha}$ are piecewise convex. Since both of them are also continuously differentiable across pieces, the conditions of Bauschke et al. [2016, Theorem 5.5] are satisfied; see also the verification argument for Bauschke et al. [2016, Example 6.1]. We thus conclude that $h+f_{\alpha}$ and $h-f_{\alpha}$ are convex on $(0,1]\times(0,1]$.
We finish the proof. ∎
Finally, we present the detailed proof of Observation 1.
Proof of Observation 1.
We set the parameter $\alpha$ and the index $k_{1}$ as follows:
$\alpha=\frac{2}{\left\lceil 40\exp\left(\frac{1}{\epsilon}\right)\right\rceil+K},\qquad k_{1}=\left\lceil 40\exp\left(\frac{1}{\epsilon}\right)\right\rceil.$
Let $k_{2}$ be the minimal index such that $x_{1}^{k}<2\alpha$, i.e., $k_{2}=\min\{k\in{\mathbb{N}}:x_{1}^{k}<2\alpha\}$. Our first goal is to prove that $k_{1}+K<k_{2}$.
A direct computation shows that $\nabla_{x_{1}}f_{\alpha}(\bm{x}^{k})>0$. Combined with the KKT conditions (9), this implies that $x_{1}^{k+1}<x_{1}^{k}\leq 1,\lambda_{1}^{k}=0$, and
$\frac{1}{x^{k+1}_{1}}=\frac{1}{x^{k}_{1}}+\frac{1}{4}\phi_{\alpha}^{\prime}(x_{1}^{k})\cdot(x_{2}^{k}+0.05).$ (22)
For the second coordinate $x_{2}$, we recall the update rule from (10):
[…]
By the definition of $k_{2}$, we have $x^{k}_{1}\geq 2\alpha$ for $k<k_{2}$. It then follows from the definition of $\phi_{\alpha}$ that $\phi_{\alpha}^{\prime}(x_{1}^{k})=2$ and $\phi_{\alpha}(x_{1}^{k})\geq 0$ for all $k<k_{2}$. Combining these with the update rule (22) and (10), we obtain $x_{2}^{k}\leq x_{2}^{0}=0.1$ and
$\frac{1}{x_{1}^{k}}<\frac{1}{x_{1}^{k-1}}+0.1<\cdots<\frac{1}{x_{1}^{0}}+0.1\cdot k=1+\frac{k}{10},\qquad\forall\,k\leq k_{2}.$ (23)
Then, armed with the fact that $x_{1}^{k_{2}}<2\alpha$ and (23), we get
$\frac{1}{x_{1}^{k_{2}}}>\frac{1}{2\alpha}\quad\text{ and }\quad\frac{1}{x_{1}^{k_{2}}}<1+\frac{k_{2}}{10}.$
Combining the two gives $k_{2}>10\cdot(\tfrac{1}{2\alpha}-1)$. Since $\alpha\leq 0.1$, it follows that
$\frac{1}{2\alpha}-1>\frac{1}{2\alpha}\cdot(1-2\alpha)\geq\frac{1}{2\alpha}\cdot(1-0.2)=\frac{0.8}{2\alpha}=\frac{4}{10\alpha}\quad\text{and }\quad k_{2}>\frac{4}{\alpha}.$
Due to our construction of $K$ and $\alpha$, we will immediately get $k_{1}+K=\frac{2}{\alpha}<k_{2}$.
Next, we are ready to continue to control $\|\bm{x}^{k}-\tilde{\bm{x}}^{\star}\|$ for $k\in[k_{1},k_{1}+K]$. Combining the update rule (10) with the fact that $\phi_{\alpha}^{\prime}(x)=2(x-\alpha)$ for $x\geq\alpha$, and using the bound $x_{1}^{k}\geq 2\alpha$ for all $k<k_{2}$, we obtain the following estimate for $x_{2}^{k}$:
$\frac{1}{x_{2}^{k}}\geq\frac{1}{x_{2}^{k-1}}+\frac{1}{2}(x_{1}^{k-1}-\alpha)\geq\frac{1}{x_{2}^{k-1}}+\frac{1}{4}x_{1}^{k-1}\geq\cdots\geq\frac{1}{x_{2}^{0}}+\frac{1}{4}\sum_{i=0}^{k-1}x_{1}^{i},\qquad\forall~{}k\leq k_{2}.$
Using $x_{2}^{0}=0.1$ and the bound from (23), i.e., $x_{1}^{i}\geq\tfrac{10}{i+10}$, we further obtain
$\frac{1}{x^{k}_{2}}\geq 10+\frac{1}{4}\sum_{i=0}^{k-1}\frac{10}{i+10}\geq 10+\frac{5}{2}\sum_{i=0}^{k-1}\int_{i}^{i+1}\frac{1}{x+10}{\rm d}x=10+\frac{5}{2}\log\left(1+\frac{k}{10}\right),\qquad\ \forall\ k\leq k_{2}.$
The bound above, together with the definition $k_{1}=\left\lceil 40\exp\left(\frac{1}{\epsilon}\right)\right\rceil$ and $k_{1}+K<k_{2}$, implies
$x_{2}^{k}\leq\left(10+\frac{5}{2}\cdot\frac{1}{\epsilon}\right)^{-1}=\frac{1}{10+\frac{5}{2}\cdot\frac{1}{\epsilon}}=\frac{2\epsilon}{20\epsilon+5}\leq\frac{2}{5}\epsilon\leq\frac{1}{2}\epsilon,\qquad\forall\,k\in[k_{1},k_{1}+K],$
where the first inequality follows from $\log(1+\tfrac{k}{10})\geq\frac{1}{\epsilon}$ for $k\geq k_{1}$.
On the other hand, using the update rule (22), along with the facts that $x_{2}^{k}>0$ and $\phi_{\alpha}^{\prime}(x^{k}_{1})=2$ for all $k<k_{2}$, we obtain
$\frac{1}{x^{k}_{1}}\geq\frac{1}{x^{k-1}_{1}}+0.025\geq\cdots\geq\frac{1}{x^{0}_{1}}+\frac{k}{40}\geq 1+\frac{k_{1}}{40}>\exp\left(\frac{1}{\epsilon}\right)>\frac{2}{\epsilon},\qquad\forall\ k\in[k_{1},k_{1}+K],$
where the last inequality holds because $\epsilon\in(0,1]$ and the function $x\mapsto\exp(x)-2x$ is positive on $[1,+\infty)$. It follows that
$x_{1}^{k}\leq\frac{\epsilon}{2},\qquad\quad\forall\ k\in[k_{1},k_{1}+K].$
Combining this with $\tilde{\bm{x}}^{\star}={\mathbf{0}}$ and the earlier bound $x_{2}^{k}\leq\tfrac{\epsilon}{2}$ for $k\in[k_{1},k_{1}+K]$, we conclude that
$\|\bm{x}^{k}-\tilde{\bm{x}}^{\star}\|\leq\epsilon,\qquad\quad\forall\ k\in[k_{1},k_{1}+K].$
This completes the proof. ∎